### v1

# DeepBooTS: Dual‑Stream Residual Boosting for Drift‑Resilient Time‑Series Forecasting

Abstract
Time-Series (TS) exhibits pronounced non-stationarity. Consequently, most forecasting methods display compromised robustness to concept drift, despite the prevalent application of instance normalization. We tackle this challenge by first analysing concept drift through a bias-variance lens and proving that weighted ensemble reduces variance without increasing bias. These insights motivate DeepBooTS, a novel end-to-end dual-stream residual-decreasing boosting method that progressively reconstructs the intrinsic signal. In our design, each block of a deep model becomes an ensemble of learners with an auxiliary output branch forming a highway to the final prediction. The block‑wise outputs correct the residuals of previous blocks, leading to a learning‑driven decomposition of both inputs and targets. This method enhances versatility and interpretability while substantially improving robustness to concept drift. Extensive experiments, including those on large-scale datasets, show that the proposed method outperforms existing methods by a large margin, yielding an average performance improvement of 15.8% across various datasets, establishing a new benchmark for TS forecasting.
Codehttps://github.com/Anoise/DeepBooTS
1 Introduction
Time Series (TS) from natural and engineered systems often evolve under transient conditions (1976_timeseries) and therefore violate stationarity assumptions (hyndman2018forecasting). Classical statistical models such as ARIMA (piccolo1990distance) and exponential smoothing (gardner1985exponential), which depend on fixed statistical properties, struggle to track these shifts (de200625). Deep neural networks have become attractive alternatives because of their powerful nonlinear modeling capabilities (hornik1991approximation), yet even the latest attention‑based (Zhou2021Informer; nie2022time; liu2023itransformer; ye2024frequency; wang2024timemixer; wang2024timexer) and graph‑neural approaches (wu2019graph; yi2024fouriergnn) provide only marginal gains over simple multilayer perceptrons and incur substantial inference overhead (shao2022spatial; zeng2023transformers; yi2024filternet; yu2025drift; yu2025learning).
Figure 1: TS forecasting models suffer from concept drift problems, more experiments are provided in Appendix H.
A deeper issue is that leading architectures like Transformers (NIPS2017_Transformer) and their variants are highly susceptible to concept drift—a mismatch between training and testing distributions. Empirical evidence shows that validation error rises early in training even as training error falls, as shown in Fig. 1, a sign that the learned model fails to generalize as the underlying data distribution shifts. Prevailing solutions, including reversible instance normalization (RevIN) (kim2021reversible) and temporal re‑aggregation, e.g., DLinear (zeng2023transformers) and iTransformer (liu2023itransformer), alleviate mean shifts but leave the variance unstable, which leads to large fluctuations and persistent concept drift. This phenomenon occurs across multiple TS forecasting models and datasets, as shown in Fig. 2(a), underscoring the need for an approach that directly targets the high‑variance component.
Figure 2: (a) Concept drift exists in current TS datasets. (b) Concept drift causes the model training error to decrease while the validation error increases. (c) The performance of DeepBooTS after reducing concept drift.
This work addresses the problem by framing concept drift through the bias-variance decomposition. We theoretically show that when the mean bias and noise are fixed, the degree of drift is controlled by the variance of the model predictions. We then prove that ensemble averaging reduces the prediction variance while preserving bias and derive new bounds showing that weighted ensembles achieve a strictly lower error under distribution shift. These results suggest that a properly designed ensemble can substantially mitigate concept drift. Building on this insight, we introduce DeepBooTS, a novel dual-stream residual-decreasing boosting architecture that progressively reinstates the intrinsic values of a non‑stationary TS. The key innovations are:
  • We rigorously analyze concept drift through the lens of the bias-variance trade-off and prove that even simple deep ensembles can substantially reduce prediction variance without increasing bias.
  • An efficient implementation of DeepBooTS is presented. Specifically, the outputs of subsequent blocks subtract the predictions of previous blocks, causing the network to explicitly model and reduce residual errors layer by layer. This residual‑learning mechanism is analogous to gradient boosting, but implemented within a deep network, enhancing robustness to distributional shifts.
  • A dual‑stream decomposition is designed that decomposes both the input and labels, enabling the model to learn complementary representations while enhancing model versatility and interpretability.
We validate DeepBooTS on a wide range of TS benchmarks, including large‑scale TS datasets. As shown in Fig. 2(c), our method consistently outperforms state‑of‑the‑art baselines, achieving an average 15.8% improvement. These results, together with our theoretical analysis, demonstrate that DeepBooTS not only advances the accuracy of TS forecasting but also provides a principled solution to the longstanding problem of concept drift. The code and datasets can be obtained from the appendices we uploaded.
2 Methodology
Figure 3: (a) Relationship among model’s bias, variance, and loss. (b) Deep boosting ensemble learning process.
2.1 Deep Ensemble helps Alleviate Concept Drift
As analyzed above, RevIN-processed (Z-Score) TS data maintains mean stationarity while exhibiting substantial variance instability. We consider a forecasting task where the data distribution shifts from an initial distribution $P_{0}$ to a new distribution $P_{t}$ at time $t$. Formally, let $Y=f_{0}(X)+\varepsilon$ under $P_{0}$ (with true regression function $f_{0}$), and at time $t$ let $Y=f_{t}(X)+\varepsilon$ under $P_{t}$, where $f_{t}$ is the new underlying function after drift, $\varepsilon\sim\mathcal{N}(0,\sigma)$. For forecasting results $\hat{Y}$, the estimation error (MSE) of the model is
$\underbrace{\text{Var}(\hat{Y})+(\text{Bias}(\hat{Y}))^{2}+\sigma^{2}}_{\text{Test Error}}=\underbrace{\mathbb{E}[(\hat{Y}-Y)^{2}]+2\mathbb{E}(\varepsilon(\hat{Y}-\mathcal{Y}))}_{\text{Training Error}}.$ (1)
[…]
Theorem 1.
Let $\hat{Y}_{1},\cdots,\hat{Y}_{N}$ be $N$ i.i.d random variables drawn from some unknown distribution, and the ensemble method is $\bar{Y}=\frac{1}{N}\sum_{t}\hat{Y}^{t}$. Then,
[…]
The proof is given in Appendix B.2. Although it adopts the simplest ensemble form, it proves that ensembling reduces variance while conserving bias. If substituting $\hat{f}_{0}(X)$ (which was trained to approximate $f_{0}$ on $P_{0}$) for $\hat{Y}$ in Eq. 1, we have
$\text{MSE}(f_{t}(X)-\hat{f}_{0}(X))=\text{Var}(\hat{f}_{0}(X))+(\text{Bias}(\hat{f}_{0}(X)))^{2}+\sigma^{2}.$ (3)
After the distribution shifts to $P_{t}$, the forecasting error is
$\text{MSE}_{P_{t}}(f_{0}(X))=\mathbb{E}_{P_{t}}[\text{Bias}_{P_{0}}(X)^{2}]+\mathbb{E}_{P_{t}}[\text{Var}_{P_{0}}(f_{0}(X))]+\sigma^{2}.$ (4)
Let $\mathrm{Var}_{P_{t}}[X]=\sigma_{t}^{2}$ with $c^{2}=\sigma^{2}_{t}/\sigma^{2}$ to quantify the change in input variance, thus we have
Theorem 2.
Given a weighted ensemble models $\hat{f}_{\text{ens}}(X)=\frac{1}{L}\sum_{l=1}^{L}\alpha_{l}g_{l}(x)$, where $\sum_{l=1}^{L}\alpha_{l}=1$, we have
$\text{MSE}_{P_{t}}(\hat{f}_{0})-\text{MSE}_{P_{t}}(\hat{f}_{\text{ens}})\approx\frac{(L-1)(1-\alpha_{l})}{L}c^{2}\sigma^{2}_{t}\geq 0.$ (5)
The proof is given in Appendix B.3. It indicates that the ensemble’s MSE under $P_{t}$ is lower than that of the single one.
2.2 Deep Ensemble Model for TS Forecasting
Based on Theorem 2, we consider a ensemble learner set $\{g_{l}\}_{l=1}^{L}$ to minimize
$\ \ \mathbb{E}(Y-\hat{f}_{\text{ens}}(X))^{2}=\mathbb{E}(Y-\sum_{l=1}^{L}\alpha_{l}g_{l}(X))^{2},$ (6)
$s.t.\ \ \sum_{l=1}^{L}\alpha_{l}=1,\text{and}\ \ \alpha_{l}\geq 0,\ \ l=1,\cdots,L.$
Here, the objective in Eq. 6 implies that
$(Y-\hat{f}_{\text{ens}}(X))^{2}=\sum_{l=1}^{L}\alpha_{l}(Y-g_{l}(X))^{2}-\sum_{l=1}^{L}\alpha_{l}(\hat{f}_{\text{ens}}(X)-g_{l}(X))^{2}.$ (7)
The proof is given in Appendix B.4. Eq. 7 proves that the ensemble generalization error can be reduced by increasing the ambiguity without increasing the bias, since the ambiguity item (the 2nd term on the right), which measures the disagreement among the base predictors, is always nonnegative. Thus, Eq. 7 indicates that we can design an ensemble learning algorithm to reduce concept drift.
Motivated by this, we propose a deep residual-decreasing ensemble method based on prediction variance minimization and TS decomposition principles, named DeepBooTS. As shown in Fig. 3(b), DeepBooTS is a two-stream architecture corresponding to Eq. 7, which is equivalent to a meta-algorithmic Boosting approach, reducing complexity of the model and thus mitigating the risk of concept drift. As shown in Fig. 3, the ‘input’ stream is obviously a decomposition of $X$ since
$X$ $=\sum_{l=0}^{L-1}g_{l}(X)+R_{L},$ (8)
where $R_{L}$ is the residual term. Further, the aim of the ‘output’ stems is to learn $L$ simple learners in a hierarchy where each learner gives more attention (larger weight) to the hard samples from the previous learner. This is equivalent to the Boosting ensemble learning process, where the final prediction is a weighted sum of $L$ simple learners, and the weights are determined by the previous learner. Let $g_{l}(x)$ denote the $l$-th learner, and $\alpha_{l}$ denote the weight of the $l$-th learner. The overall estimation $\hat{f}_{\text{ens}}(X)$ is a weighted subtraction of the $L$ estimations. For a sample $X$, we have
$\hat{f}_{\text{ens}}(X)=i\sum_{l=0}^{\hbar}\alpha_{2l+1}g_{2l+1}(X)-i\sum_{l=0}^{\hbar}\alpha_{2l}g_{2l}(X),$ (9)
$s.t.\ \ i\sum_{l=0}^{\hbar}\alpha_{2l+1}-i\sum_{l=0}^{\hbar}\alpha_{2l}=1,\text{and}\ \ \alpha_{l}\geq 0.$
where $i=1$ if $L\bmod 2=1$, else i=1i=-1, and $\hbar=\lfloor\frac{L}{2}\rfloor$. Similar to Eq 6, Eq. 9 can be rewritten as
$(Y-\hat{f}_{\text{ens}}(X))^{2}$ $=\sum_{l=0}^{\hbar}i(\alpha_{2l+1}-\alpha_{2l})(Y-g_{l}(X))^{2}$
$-\sum_{l=0}^{\hbar}i(\alpha_{2l+1}-\alpha_{2l})(\hat{f}_{\text{ens}}(X)-g_{l}(X))^{2}$ (10)
Obviously, Eq. 2.2 derived from Eq. 9 is similar to Eq. 7. However, Eq. 9 has more advantages, that is
Theorem 3.
Assume that the estimation error of block $g_{l}(X)$ is $e_{l}$, $e_{l}\overset{i.i.d}{\sim}\mathcal{N}(0,\nu)$. Let $\alpha_{l}=\alpha$ be the weight of $g_{l}$, $l\in[0,L]$, and the covariance of estimations of two different blocks by $\mu$, we have
$\text{Var}(\hat{Y})<\frac{4}{L}\alpha^{2}(\nu+\mu).$ (11)
The proof is given in Appendix B.5. Clearly, the variance of DeepBooTS is bound by the estimation error (noise error) of each block, and the covariance between blocks. It is evident that the subtraction adopted in DeepBooTS can reduce the variance, thereby mitigating concept drift. On the contrary, switching the aggregation operation of the output stream in the DeepBooTS to addition results in an approximate variance of $\frac{4}{L}\alpha^{2}\nu+3\alpha^{2}\mu$, which is much larger than that of subtraction in Eq. 9. This phenomenon is also supported by our experiments in Section 3.5. Furthermore, Theorem 3 also demonstrates that increasing the number of layers $L$ does not escalate the risk of concept drift. It proves that the deep ensemble models can go deeper, and the test error approaches $\text{Bias}(\hat{f}_{\text{ens}}(X))^{2}+\sigma^{2}$ when $L$ is infinite and doesn’t consider performance-efficient trade-off.
2.3 Implementation
Here, we employ deep neural networks to implement DeepBooTS. As shown in Fig. 4, the model consists of two primary data streams. One is the input stream decomposed through multiple residual learners, while the other is the output stream that progressively learns the residuals of the supervised signals. Along the way, they pass through multiple learners capable of converting signals. The model is simple and versatile, yet powerful and interpretable. The pseudocode is given in Appendix L, and complexity analysis is given in Appendix G. We now delve into how these properties are incorporated into DeepBooTS.
Backbone: The fundamental building backbone features a fork architecture, which accepts one input $X_{l}$ and produces two distinct streams, $R_{l}$ and $O_{l}$. Concretely, $R_{l}$ is the remaining portion of $X_{l}$ after it has undergone processing within a neural module, which can be expressed as
$\hat{X}_{l}=\text{Block}_{l}(X_{l}),\qquad R_{l}=X_{l}-\hat{X}_{l},$ (12)
Equation 12 represents an implicit decomposition of $X$, which differs from the moving average adopted by (wu2021autoformer; zhou2022fedformer; liang2023does) but is similar to (oreshkin2019n). Differently, the residual $R_{l}$ captures what remains unaltered, providing a basis for comparison with the transformed portion.
In the subsequent steps, the intention is to maximize the utilization of the subtracted portion $\hat{X}_{l}$. First, $\hat{X}_{l}$ is projected into the same dimension as the label that is anticipated, $Y$. This process can be expressed as
$O_{l}$ $=\text{Predictor}_{l}(\hat{X}_{l}),$ (13)
where $O_{l}$ is the prediction results of the $l$-th predictor. Then, $O_{l}$ will be subtracted from the outputs of the next predictor sequentially until the final prediction $\hat{Y}$ is achieved.
Learner: The basic learner can be constructed utilizing widely-used neural structures, such as fully connected, convolutional, and attention layers. A ready-made solution is to utilize Attention to learn subtle relationships between attributes. As analyzed in Section 2.2, we implement a corrective measure by subtracting the output from the input, i.e.,
$\hat{X}_{l,1}=\text{Attention}_{\theta_{l,1}}(X_{l,1}),\quad R_{l,1}=X_{l,1}-\delta\hat{X}_{l,1},$ (14)
where $\delta$ is 1 when the module exists, otherwise it is 0. $\delta$ serves to eliminate the Attention layer when it exerts adverse effects, thereby enabling the unimpeded flow of input towards the FeedForward (FF) layers:
$\hat{X}_{l,2}=\text{FF}_{\theta_{l,2}}(R_{l,1}),\qquad R_{l,2}=X_{l,2}-\hat{X}_{l,2}.$ (15)
Also, the Attention can be implemented in the frequency domain by Fast Fourier Transform (FFT) to achieve lightweight and efficient processing, as
$Q_{l},K_{l},V_{l}=\text{FFT}_{\theta_{Q},\theta_{K},\theta_{V}}(X_{l}),$ (16)
$\hat{X}_{l,1}=\text{FFT}^{-1}(\phi(Q_{l}K_{l}^{T})V).$ (17)
where $\phi$ is the activation function.
Figure 4: The implementation of DeepBooTS.
Coefficient: Upon being scaled by a coefficient ($\alpha_{l}$), the dual streams are directed toward the next block or projected into the output space. Here, we adopt a learnable gating mechanism as the coefficient. The aim is to let each learner regulate the stream-transmission pace autonomously. Thus, for input and output streams, we have
$X_{l+1}$ $=\varphi(\theta_{1}(R_{l,2}))\cdot\theta_{2}(R_{l,2}),$ (18)
${O}_{l+1}$ $=\varphi(\theta_{3}([\hat{X}_{l,1},\hat{X}_{l,2}]))\cdot\theta_{4}([\hat{X}_{l,1},\hat{X}_{l,2}]),$ (19)
where $\theta_{1}$ and $\theta_{2}$ are learnable parameters, $\varphi$ is the sigmoid, and the brackets ‘[ ]’ are a concatenation operation.
Model DeepBooTS* DeepBooTS TimeXer TimeMixer iTransformer PatchTST Crossformer DLinear FEDformer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETT 0.346 0.378 0.362 0.382 0.365 0.388 0.367 0.388 0.383 0.407 0.381 0.397 0.942 0.684 0.559 0.515 0.437 0.449
Traffic 0.373 0.261 0.406 0.270 0.466 0.287 0.484 0.297 0.428 0.282 0.481 0.304 0.550 0.304 0.625 0.383 0.610 0.376
ELC 0.158 0.252 0.166 0.259 0.171 0.270 0.182 0.272 0.178 0.270 0.205 0.290 0.244 0.334 0.212 0.300 0.214 0.327
Weather 0.227 0.266 0.245 0.271 0.241 0.271 0.240 0.271 0.258 0.279 0.259 0.281 0.259 0.315 0.265 0.317 0.309 0.360
Solar 0.197 0.240 0.227 0.250 0.272 0.300 0.216 0.280 0.233 0.262 0.270 0.307 0.641 0.639 0.330 0.401 0.291 0.381
PEMS 0.075 0.178 0.109 0.218 0.141 0.225 0.138 0.242 0.113 0.221 0.180 0.291 0.169 0.281 0.278 0.375 0.213 0.327
$1^{st}$ or $2^{st}$ 30 30 15 22 3 4 8 2 3 2 0 0 3 0 1 0 0 0
Table 1: Multivariate TS forecasting results (average). The results for all forecasting lengths are provided in Appendix M.
Model DeepBooTS Periodformer FEDformer Autoformer Informer LogTrans Reformer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.072 0.206 0.093 0.237 0.111 0.257 0.105 0.252 0.199 0.377 0.345 0.513 0.624 0.600
ETTh2 0.185 0.337 0.192 0.343 0.206 0.350 0.218 0.364 0.243 0.400 0.252 0.408 3.472 1.283
ETTm1 0.052 0.172 0.059 0.201 0.069 0.202 0.081 0.221 0.281 0.441 0.231 0.382 0.523 0.536
ETTm2 0.118 0.254 0.115 0.253 0.135 0.278 0.130 0.271 0.147 0.293 0.130 0.277 0.136 0.288
Traffic 0.132 0.212 0.150 0.233 0.177 0.270 0.261 0.365 0.309 0.388 0.341 0.417 0.375 0.434
Electricity 0.314 0.401 0.298 0.389 0.347 0.434 0.414 0.479 0.372 0.444 0.410 0.473 0.352 0.435
Weather 0.0015 0.0293 0.0017 0.0317 0.008 0.067 0.0083 0.0700 0.0033 0.0438 0.0059 0.0563 0.0115 0.0785
Exchange 0.429 0.453 0.353 0.434 0.499 0.512 0.578 0.537 1.511 1.029 1.350 0.810 1.028 0.812
$1^{\text{st}}$ Count 24 27 14 12 0 1 0 0 0 0 2 0 0 0
Table 2: Univariate TS forecasting results (average). The results for all forecasting lengths are provided in Appendix N.
3 Experiments
We conduct a comprehensive comparison with the 18 latest SOTA methods, including TimeXer (wang2024timexer), TimeMixer (wang2024timemixer), iTransformer (liu2023itransformer), PatchTST (nie2022time), Crossformer (zhang2022crossformer), SCINet (liu2022scinet), DLinear (zeng2023transformers), FEDformer (zhou2022fedformer) for multivariate forecasting, and FreTS (yi2024frequency), PSLD (liang2024psld), and FourierGNN (yi2024fouriergnn) for large-scale TS datasets, as well as Periodformer (liang2023does), Autoformer (wu2021autoformer), Informer (Zhou2021Informer), LogTrans (li2019enhancing), Reformer (Kitaev2020Reformer), N-BEATS (oreshkin2019n) and N-Hits (challu2023nhits) for univariate forecasting. More experimental settings can be found in Appendix D.
3.1 Main Experimental Results
All datasets are adopted for both multivariate (multivariate predict multivariate) and univariate (univariate predicts univariate) tasks. The detailed information pertaining to the datasets can be located in Appendix C. The models used in the experiments are evaluated over a wide range of prediction lengths to compare performance on different future horizons: 96, 192, 336 and 720. Please refer to Appendix K for more experiments on the full ETT dataset.
Multivariate Results: The results for multivariate TS forecasting are outlined in Table 1, with the optimal results highlighted in red and the second-best results emphasized with underlined. Due to variations in input lengths among different methods, for instance, PatchTST and DLinear employing an input length of 336, while Crossformer and Periodformer search for the input length without surpassing the maximum setting (720 in Crossformer, 144 in Periodformer), we have configured two versions of DeepBooTS: 336 for DeepBooTS* and 96 for DeepBooTS.
As shown in Table 1, DeepBooTS achieves the consistent SOTA performance across all datasets and prediction length configurations. Compared with the latest models acknowledged for their exceptional average performance, the proposed DeepBooTS demonstrates an average performance increase of 15.8%, achieving a substantial performance improvement. Obviously, it achieves advanced performance, averaging 30 items across six datasets, with an improvement in average performance on each dataset. These experimental results confirm that the proposed DeepBooTS demonstrates superior prediction performance across different datasets with varying horizons. For a more detailed analysis, please refer to Appendix O.
[…]
3.2 Evaluation on Monash TS Datasets
Further, we evaluate the proposed method on 7 Monash TS datasets (godahewa2021monash) (e.g., NN5, M4 and Sunspot, etc.) and 7 diverse metrics (e.g., MAPE, sMAPE, MASE and Quantile, etc.) to systematically evaluate our model. All experiments are compared under the same input length (e.g., II=96) and output lengths (e.g., OO={96, 192, 336 and 720}). As shown in Table 3, the proposed DeepBooTS emerged as the frontrunner, achieving a score of 41 out of 54. Please refer to Appendices Q and R for details about the definition and experimental settings, respectively.
Mulvariate ILI Oik_Weather NN5
Metric MSE MAE RMSP sMAPE MASE Q75 MSE MAE RMSP sMAPE MASE Q75 MSE MAE RMSP sMAPE MASE Q75
DeepBooTS 2.016 0.86 0.367 0.650 0.574 0.955 0.689 0.618 0.773 1.082 0.888 0.628 0.719 0.578 0.541 0.861 0.530 0.559
iTransformer 2.262 0.958 0.415 0.703 0.648 1.107 0.720 0.632 0.781 1.090 0.935 0.636 0.723 0.583 0.554 0.877 0.542 0.565
DLinear 2.915 1.188 0.574 0.865 0.791 1.524 0.701 0.637 0.805 1.185 0.943 0.641 1.448 0.929 0.969 1.304 0.835 0.978
Autoformer 3.187 1.224 0.596 0.909 0.740 1.349 0.853 0.712 0.883 1.201 0.992 0.704 0.851 0.659 0.636 0.975 0.603 0.637
Informer 5.208 1.576 0.792 1.183 0.907 2.304 0.67 0.625 0.795 1.126 0.818 0.640 0.967 0.727 0.739 1.081 0.662 0.722
Univariate M4 Hourly Us_births Saugeenday
DeepBooTS 0.223 0.287 0.230 0.509 0.276 0.311 0.364 0.431 0.248 0.445 0.363 0.361 1.194 0.544 0.808 1.053 1.232 0.674
iTransformer 0.310 0.390 0.376 0.656 0.366 0.410 0.378 0.439 0.251 0.438 0.365 0.388 1.206 0.552 0.832 1.066 1.258 0.687
N-Beats 0.337 0.423 0.392 0.700 0.390 0.493 0.670 0.638 0.409 0.625 0.546 0.834 1.028 0.543 0.962 1.387 0.943 0.572
N-Hits 0.346 0.418 0.402 0.687 0.396 0.442 0.538 0.578 0.364 0.540 0.464 0.802 0.982 0.532 0.865 1.189 0.858 0.562
Autoformer 0.635 0.611 0.614 0.881 0.518 0.741 0.901 0.757 0.447 0.789 0.737 0.788 1.252 0.640 1.058 1.370 1.057 0.674
    All results are averaged across all prediction lengths. The results for all prediction lengths and the experimental settings are provided in Appendix R. The definitions of all metrics are provided in Appendix Q.
Table 3: Mulvariate and univariate forecasting results with diverse metrics on Monash TS datasets.
Figure 5: (a) Comparisons on large-scale TS datasets. (b) Visualization depicting the output of each block in DeepBooTS.
3.3 Performance on Large-Scale Datasets
The performance comparisons for large-scale TS datasets are summarized in Fig. 5(a), including the CBS dataset with 4,454 nodes (17GB) and the Milano dataset with 10,000 nodes (19GB). On these datasets, it is not feasible to place all nodes on a single GPU for long-term TS forecasting. Therefore, we adopted the random partitioning strategy (liang2024act) for our experiments. Please refer to Appendices C.3 and D for details. Compared to the latest advanced PSLD, the proposed DeepBooTS yields an overall MSE reduction of 8.9% and 6.2% on the CBS and Milano datasets, respectively. In addition, we are pleased to report that our method has been successfully applied to real-world scenarios, e.g., energy and power, showing substantial practical value. For more experiments, please refer to Appendix F.
Figure 6: Ablation studies of DeepBooTS using various Attention. All results are averaged across all prediction lengths. The tick labels of the X-axis are the abbreviation of Attention types. The detailed setup and results are provided in Appendix J.
3.4 Generality
[…]
3.5 Effectiveness
Figure 7: Ablation studies on various components of DeepBooTS. All results are averaged across all prediction lengths. The variables $X$ and $Y$ represent the input and output streams, while the signs ‘+’ and ‘-’ denote the addition or subtraction operations used when the streams’ aggregation. The letter ‘G’ denotes adding a learnable gating (Coefficient) to the output of each block.
[…]
3.6 Interpretability
The intrinsic characteristic of DeepBooTS lies in the alignment of the output from each block with the final output. This alignment, in turn, expedites the interpretability of the model’s learning process. As depicted in Fig. 5(b), the output of each block in DeepBooTS is visualized. It becomes evident that each block discerns and assimilates meaningful patterns. Specifically, when the embedding dimension is low, each block must learn salient patterns. Further, when the depth of the model is increased, it becomes evident that the amplitude of the shallow block decreases, and numerous components are transferred to the deep block. For a more detailed analysis, please refer to Appendix P.
3.7 Reduce Variance and Go Deeper
Figure 8: Comparisons of model’s variance and depth.
By adjusting model hyperparameters and injecting noise, we generated a diverse set of predictions from latest advanced models. Then, these predictions are statistically analyzed to compute the prediction mean and variance. As shown in Fig. 8(a), DeepBooTS achieves superior performance and the smallest prediction variance, while other models exhibit weaker alignment with ground truth and higher variance. This validates the effectiveness of our theory and model design. Meanwhile, these results also confirm that DeepBooTS is less sensitive to hyperparameters, with details presented in Appendix I. Furthermore, given DeepBooTS’s robustness against high-variance, it can be designed with considerable depth. Fig. 8(b) shows the scenarios when models go deeper. Serious overfitting happens when the number of iTransformer blocks is increased from 4 to 8. However, even with the DeepBooTS blocks deepened to 16, it continues to exhibit excellent performance.
4 Conclusion
In this paper, we theoretically analyze the impact of concept drift in TS forecasting from a bias-variance perspective. We demonstrate that the deep ensemble is capable of reducing this issue. Based on this, we propose a novel dual-stream residual-decreasing Boosting ensemble approach, termed DeepBooTS, and demonstrate its efficacy in variance reduction. DeepBooTS facilitates the learning-driven and progressive decomposition of both the input and output streams, thereby empowering the model resilience against concept drift. Extensive experiments show that DeepBooTS achieves SOTA performance, and it is less sensitive to hyperparameters and can be designed very deep with enhanced interpretability. Furthermore, DeepBooTS can serve as a versatile framework to enhance other models’ performance.
5 Acknowledgments
This work was supported by Project of Key R&D Program of Shandong Province, China (2025CXGC010107, 2025CXPT095), Taishan Scholars Program: NO.tspd20240814, the Pilot Project for the Integration of Science, Education and Industry of Qilu University of Technology (Shandong Academy of Sciences)(2025ZDZX01), National K&D Development Program of China (2024YFB3312302), and the Qilu Institute of Technology for Talent Project (QIT24TP027), and the Young Talent of Lifting engineering for Science and Technology in Shandong (SDAST2025QTB008).
Appendix A Related Work
A.1 Classical Models for TS Forecasting
TS forecasting is a classic research field where numerous methods have been invented to utilize historical series to predict future missing values. Early classical methods (piccolo1990distance; gardner1985exponential) are widely applied because of their well-defined theoretical guarantee and interpretability. For example, ARIMA (piccolo1990distance) initially transforms a non-stationary TS into a stationary one via differencing, and subsequently approximates it using a linear model with several parameters. Exponential smoothing (gardner1985exponential) predicts outcomes at future horizons by computing a weighted average across historical data. In addition, some regression-based methods, e.g., random forest regression (RFR) (liaw2002classification) and support vector regression (SVR) (castro2009online), etc., are also applied to TS forecasting. These methods are straightforward and have fewer parameters to tune, making them a reliable workhorse for TS forecasting. However, their shortcoming is insufficient data fitting ability, especially for high-dimensional series, resulting in limited performance.
A.2 Deep Models for TS Forecasting
The advancement of deep learning has greatly boosted the progress of TS forecasting. Specifically, convolutional neural networks (CNNs) (lecun1998gradient) and recurrent neural networks (RNNs) (connor1994recurrent) have been adopted by many works to model nonlinear dependencies of TS, e.g., LSTNet (lai2018modeling) improve CNNs by adding recursive skip connections to capture long- and short-term temporal patterns; DeepAR (salinas2020deepar) predicts the probability distribution by combining autoregressive methods and RNNs. Several works have improved the series aggregation forms of Attention mechanism, such as operations of exponential intervals adopted in LogTrans (li2019enhancing), ProbSparse activations in Informer (Zhou2021Informer), frequency sampling in FEDformer (zhou2022fedformer) and iterative refinement in Scaleformer (shabani2022scaleformer). Besides, GNNs and Temporal convolutional networks (TCNs) (lea2016temporal) have been utilized in some methods (wu2019graph; li2023dynamic; liu2022scinet; wu2022timesnet) for TS forecasting on graph data. The aforementioned methods solely concentrate on the forms of aggregating input series, overlooking the challenges posed by the concept drift problem.
Since TS exhibit a variety of patterns, it is meaningful and beneficial to decompose them into several components, each representing an underlying category of patterns that evolving over time (1976_timeseries). Several methods, e.g., STL (cleveland1990stl), Prophet (taylor2018forecasting) and N-BEATS (oreshkin2019n), commonly utilize decomposition as a preprocessing phase on historical series. There are also some methods, e.g., Autoformer (wu2021autoformer), FEDformer (zhou2022fedformer), Non-stationary Transformers (liu2022non) and DistPred (liang2025distpred), that harness decomposition into the Attention module. The aforementioned methods attempt to apply decomposition to input series to enhance predictability, reduce computational complexity, or ameliorate the adverse effects of non-stationarity. Nevertheless, these prevalent methods are susceptible to significant concept drift when applied to non-stationary TS.
Furthermore, there are four themes that use deep learning to predict time series: (1) smarter transformers (NIPS2017_Transformer), such as PatchTST (nie2022time), iTransformer (liu2023itransformer) BasisFormer (ni2023basisformer), and TimeXer (wang2024timexer) which restructure attention or add learnable bases to extend context length, cut computation and boost accuracy; (2) competitive non-transformer backbones, including N-HiTS (hierarchical MLP) (challu2023nhits), DLinear (zeng2023transformers), PGN (jia2024pgn) and state-space models like TSMamba (ma2024mamba), TimeMachine (ahamed2024timemachine) and FLDMamba (zhang2025fldmamba), which deliver linear-time inference and rival or surpass transformers on long horizons; (3) foundation-model initiatives, TimeGPT (garza2023timegpt), OneFitAll (zhou2023one), TimeLLM (jin2024timellm), UniTime (liu2024unitime) and DAM (darlowdam2024DAM) that pre-train on massive heterogeneous corpora and achieve impressive zero-shot or few-shot performance across domains; and (4) training and interpretability advances, such as frequency-adaptive normalization (FAN) (ye2024frequency), e.g., FreTS (yi2024frequency), FilterNet (yi2024filternet), and decomposition-aware architectures, e.g., which tackle non-stationarity, quantify uncertainty and make forecasts more transparent.
A.3 Concept Drift in TS forecasting
Concept drift in TS forecasting has long been managed with adaptive retraining and change detection techniques.
Classical methods include using sliding windows of recent data to regularly update models, thereby “forgetting” outdated patterns. Traditional statistical forecasters like ARIMA (piccolo1990distance) can be incrementally retrained on rolling windows, e.g., an incremental ARIMA scheme fits a new model whenever the current model’s validation error exceeds a threshold, signaling a drift. Another staple is change-point detection, where statistical tests (e.g. the Page-Hinkley test) monitor the data or error distribution and flag abrupt shifts once differences exceed a predefined threshold. For instance, the Error Intersection Approach (EIA) (baier2020handling) employs two models, one “stable” (more complex) and one “reactive” (simpler), and compares their performance. Furthermore, the Drift Detection Method (DDM) (korycki2023adversarial) monitors the classifier’s error rate distribution and raises an alarm when the error’s increase is statistically significant. These classical and statistical methods form the foundation of drift handling in forecasting, using either explicit retraining policies or specialized detectors to maintain model validity over time.
Online learning strategies embed concept drift adaptation within the forecasting models themselves. One example is FSNet (pham2023learning), which leverages complementary learning systems theory to pair a slow-learning base forecaster with fast-adapting components. Another line of work is OneNet (wen2023onenet), an online ensembling approach that dynamically combines two neural models: one specializes in capturing temporal dependencies within each series, and the other focuses on cross-series (covariate) relationships. Each of these deep learning techniques illustrates how integrating drift-awareness (through dual-model architectures, ensembling, or proactive adjustment) can improve TS forecasting performance in online.
Appendix B Theoretical Rationale
B.1 Proof of Equation 1
In this subsection, we provide a theoretical analysis of how DeepBooTS alleviates concept drift. In practice, observations $Y$ often contain additive noise $\varepsilon$: $Y=\mathcal{Y}+\varepsilon$, where $\varepsilon\sim\mathcal{N}(0,\sigma)$. Then, the estimation error (MSE) of the final model is:
$\mathbb{E}[(\hat{Y}-Y)^{2}]$ $=\mathbb{E}[(\hat{Y}-\mathcal{Y}-\varepsilon)^{2}]$
$=\mathbb{E}[(\hat{Y}-\mathcal{Y})^{2}]+\mathbb{E}(\varepsilon^{2})-2\mathbb{E}(\varepsilon(\hat{Y}-\mathcal{Y}))$
$=\mathbb{E}\left((\hat{Y}-\mathbb{E}(\hat{Y})+\mathbb{E}(\hat{Y})-\mathcal{Y})^{2}\right)+\sigma^{2}-2\mathbb{E}(\varepsilon(\hat{Y}-\mathcal{Y}))$
$=\mathbb{E}\left((\hat{Y}-\mathbb{E}(\hat{Y}))^{2}+(\mathbb{E}(\hat{Y})-\mathcal{Y})^{2}+2(\hat{Y}-\mathbb{E}(\hat{Y}))(\mathbb{E}(\hat{Y})-\mathcal{Y})\right)+\sigma^{2}-2\mathbb{E}(\varepsilon(\hat{Y}-\mathcal{Y}))$
$=\mathbb{E}\left((\hat{Y}-\mathbb{E}(\hat{Y}))^{2}\right)+(\mathbb{E}(\hat{Y})-\mathcal{Y})^{2}+\sigma^{2}-2\mathbb{E}(\varepsilon(\hat{Y}-\mathcal{Y}))$
$=\text{Var}(\hat{Y})+(\text{Bias}(\hat{Y}))^{2}+\sigma^{2}-2\mathbb{E}(\varepsilon(\hat{Y}-\mathcal{Y})).$ (20)
Based on this, Equation 20 can be reformulated as
$\underbrace{\text{Var}(\hat{Y})+(\text{Bias}(\hat{Y}))^{2}+\sigma^{2}}_{\text{Test Error}}=\underbrace{\mathbb{E}[(\hat{Y}-Y)^{2}]+2\mathbb{E}(\varepsilon(\hat{Y}-\mathcal{Y}))}_{\text{Training Error}}.$ (21)
Equation 21 demonstrates a triangular relationship between the loss, bias, and variance of the model when the noise term is not taken into account, as shown in Fig. 3. For modern complex deep learning models, their biases are typically very low (goodfellow2016deep; zhang2021understanding). Here, we focus on how to reduce their variance to alleviate concept drift. In Equation 9, the final model is a weighted sum of $L$ simple models, with the weights determined by the performance of the previous model.
B.2 Proof of Theorem 1
Before we prove Theorem 1, we first give several definitions and prove a lemma.
Mean Squared Error (MSE), MSE is denoted as
$\mathcal{L}(Y,\hat{Y})=\frac{1}{2}(Y-\hat{Y})^{2}.$ (22)
[…]
Convex Conjugate: The convex conjugate of a function $g(X)$ is defined as
[…]
Central Prediction: The central prediction is the unique minimizer, i.e.,
$\mathbb{C}_{Z}=\mathop{\text{argmin}}\limits_{Z}\mathbb{E}[\mathcal{L}(\hat{Y},Z)],$ (24)
According to the above definition, we have
Lemma 4.
Let $Y_{1},\cdots,Y_{N}$ be $N$ i.i.d random variables drawn from some unknown distribution, for average ensemble $\bar{Y}=\frac{1}{N}\sum_{t}\hat{Y}^{t}$ with MSE metric, we have $\mathbb{C}_{\bar{Y}}=\mathbb{C}_{\hat{Y}}$.
Proof.
First, the conjugate function of MSE is
[…]
$=\mathop{\text{sup}}\limits_{Z}\left(ZX-\frac{1}{2}(Y-\hat{Y})^{2}\right)$
$=\mathop{\text{sup}}\limits_{Z}\left(ZX-\frac{1}{2}Y^{2}+Y\hat{Y}-\frac{1}{2}\hat{Y}^{2}\right)$
$=\mathop{\text{sup}}\limits_{Z}\left(-\frac{1}{2}Y^{2}+(Z+\hat{Y})Y-\frac{1}{2}\hat{Y}^{2}\right)$ (25)
The Eq. 25 is a quadratic function in $Y$, which opens downward (coefficient of $Y^{2}$ is negative). The maximum occurs at the vertex of the parabola, given by $Y=Z+\hat{Y}$. Thus, Substitute it back into the quadratic expression in Eq. 25, we have
$g^{*}(Z)$ $=-\frac{1}{2}(Z+\hat{Y})^{2}+(Z+2\hat{Y})(Z+2\hat{Y})-\frac{1}{2}\hat{Y}^{2}$
$=\frac{1}{2}((Z+\hat{Y})^{2}-\hat{Y}^{2})$
$=\frac{1}{2}Z^{2}-\hat{Y}Z$ (26)
On the one hand, for $\mathbb{C}_{\hat{Y}}$, we have
$\mathbb{C}_{\hat{Y}}$ $=\mathop{\text{argmin}}\limits_{Z}\mathbb{E}(\hat{Y}-Z)^{2}$
$=\mathop{\text{argmin}}\limits_{Z}\left[Z^{2}-2Z\mathbb{E}[\hat{Y}]+\mathbb{E}[\hat{Y}^{2}]\right]$ (27)
This is a quadratic function of $\hat{Y}$, and its minimum occurs at the vertex, we have
$Z=\mathbb{E}[\hat{Y}],\qquad\text{and}\qquad\mathbb{C}_{\hat{Y}}=\mathbb{E}[\hat{Y}$]. (28)
On the other hand, Substituting $\bar{Y}=\frac{1}{N}\sum_{t}\hat{Y}^{t}$ into Eq. 24, we compute $\mathbb{E}[\bar{Y}]$ using the same derivation in Eq. 28, we have $\mathbb{C}_{\bar{Y}}=\mathbb{E}[\bar{Y}]=\frac{1}{N}\sum_{t}\hat{Y}^{t}=\mathbb{E}[\hat{Y}$]. Therefore, $\mathbb{C}_{\bar{Y}}=\mathbb{C}_{\hat{Y}}$ is held. ∎
Next, we prove Theorem 1. Theorem 1 states that: Let $\hat{Y}_{1},\cdots,\hat{Y}_{N}$ be $N$ i.i.d random variables drawn from some unknown distribution, and the ensemble method is $\bar{Y}=\frac{1}{N}\sum_{t}\hat{Y}^{t})$. Using MSE as the metric, we have,
[…]
Proof.
[…]
$=\mathbb{E}[\mathcal{L}(\frac{1}{N}\sum_{i}\hat{Y_{i}}^{*},\mathbb{C}_{\hat{Y}}^{*})]$
$\leq\frac{1}{N}\sum_{i}\mathbb{E}[\mathcal{L}(\hat{Y}_{i}^{*},\mathbb{C}_{\hat{Y}}^{*})]\qquad(\text{by convexity of}\ \mathcal{L})$
$\leq\frac{1}{N}\sum_{i}\mathbb{E}[\mathcal{L}(\mathbb{C}_{\hat{Y}},\hat{Y}_{i})]$
[…]
Thus, the theorem is proved. ∎
B.3 Proof of Theorem 2
Proof.
Since the model normalizes (RevIN) by the instance’s mean and standard deviation (std) before prediction, then rescales the prediction by the instance’s original mean and std. Formally, if $Z=(X-\mu_{X})/\sigma_{X}$ is the normalized input (with $\mu_{X},\sigma_{X}$ the instance mean and std), and $h(Z)$ is the model’s prediction in normalized space, then the final prediction is:
$\hat{Y}=\mu_{X}^{Y}+\sigma_{X}^{Y}\,h(Z),$ (32)
where $\mu_{X}^{Y},\sigma_{X}^{Y}$ are the output rescaling. If an input’s deviation from the mean is scaled by factor $c$, the model’s output deviations scale by approximately the same factor $c$ due to RevIN. Thus the prediction variance changes with input variance.
Single-Model Bias and Variance Under Covariate Shift
Since the true function $f_{0}(X)$ is unchanged and models are not updated after drift, the model was approximately unbiased for $f_{0}$ on $P_{0}$, it remains so on $P_{t}$. Thus, for each model $\hat{f}$, we assume $\text{Bias}(X)\approx 0$ or small for all $X$. Therefore, we focus on variance differences, noting $E_{P_{t}}[\text{Bias}(X)^{2}]\approx E_{P_{0}}[\text{Bias}(X)^{2}]$. Now let’s observe the change of variance. The predictive variance of model $\hat{f}$ changes under $P_{t}$. Consider the variance of $\hat{f}(X)$ when $X$ is drawn from $P_{0}$. Formally,
$\text{Var}_{P_{t}}[\hat{f}(X)]=E_{X\sim P_{t}}\big[(\hat{f}(X)-E_{P_{t}}[\hat{f}(X)])^{2}\big].$ (33)
Thus, one can expect $\text{Var}_{P_{t}}[\hat{f}(X)]\approx c^{2},\text{Var}_{P_{0}}[\hat{f}(X)]$. In other words, if $\sigma_{t}^{2}=c^{2}\sigma^{2}$, then the model’s output variance scales by about $c^{2}$. Thus the expected variance term is
[…]
Therefore, for single model, $\text{Bias}^{2}_{P_{t}}\approx\text{Bias}^{2}_{P_{0}}$, $\text{Var}_{P_{t}}\approx c^{2}\text{Var}_{P_{0}}=c^{2}\sigma^{2}$, but the noise item is the same.
Ensemble Variance Reduction under $P_{t}$
Now consider the ensemble predictor $\hat{f}_{\text{ens}}(X)=\frac{1}{L}\sum_{l=1}^{L}g_{i}(X)$. The ensemble bias is $\text{Bias}_{\text{ens}}(X)=E_{P_{t}}[\hat{f}_{\text{ens}}(X)]-\hat{f}_{0}(X)=\frac{1}{L}\sum_{i=1}^{L}E_{P_{t}}[g_{l}(X)]-\hat{f}_{0}(X)=\frac{1}{L}\sum_{i=1}^{L}(g_{l}(X)+\text{Bias}(X))-\hat{f}_{0}(X)$. Since each model is unbiased ($\text{Bias}_{\text{ens}}(X)=0$), then clearly $\text{Bias}_{\text{ens}}(X)=0$. In the worst case that all models have the same bias, the ensemble’s bias is the same $\text{Bias}_{\text{ens}}(X)$. Further, the variance of the average is $\frac{1}{L^{2}}$ times the sum of their variances
$\text{Var}_{{P_{t}}}\big(\hat{f}_{\text{ens}}(X)\big)=\text{Var}_{P_{t}}\Big(\frac{1}{L}\sum_{i=1}^{L}g_{l}(X)\Big)=\frac{1}{L^{2}}\sum_{i,j}\text{Cov}(g_{i}(X),g_{j}(X)).$ (35)
If model errors are uncorrelated, $\text{Cov}(g_{l},g_{j})=0$ for $l\neq j$, and $\text{Var}(g_{l})$ is the same $\forall l$. Then this simplifies to
$\text{Var}_{P_{t}}\big(\hat{f}_{\text{ens}}(X)\big)=\frac{1}{L^{2}}\Big(L\text{Var}_{P_{t}}(g_{l}(X))\Big)=\frac{1}{L}\text{Var}_{P_{t}}(g_{l}(X)).$ (36)
In the general case, let $\alpha_{l}\in(0,1)$ is the average pairwise correlation between model prediction errors, one can show
$\text{Var}_{P_{t}}(\hat{f}_{\text{ens}}(X))=\frac{1}{L^{2}}\Big(Lc^{2}\sigma^{2}+L(L-1)\alpha_{l}\sigma^{2}\Big)=\frac{1+(L-1)\alpha_{l}}{L}c^{2}\sigma^{2},$ (37)
Thus, ensembling $L$ predictors cuts down the variance roughly by a factor of $L$, while the ensemble bias term $\text{Bias}_{\text{ens}}(X)$ is not worse than the single-model bias.
Finally, we compare the overall prediction error of a single model to that of the ensemble on the shifted distribution $P_{t}$. Using the bias–variance decomposition as shown in Eq. 4, we have
$\text{MSE}_{P_{t}}(\hat{f}_{0})-\text{MSE}_{P_{t}}(\hat{f}_{\text{ens}})\approx c^{2}\sigma^{2}-\frac{1+(L-1)\alpha_{l}}{L}c^{2}\sigma^{2}=\frac{(L-1)(1-\alpha_{l})}{L}c^{2}\sigma^{2}_{t}\geq 0.$ (38)
Therefore, under $P_{t}$, the ensemble’s total prediction error remains lower than that of any single model. ∎
B.4 Proof of Equation 6
Here, we want to prove that Eq. 6 yields
$(Y-\hat{f}_{\text{ens}}(X))^{2}=\sum_{l=1}^{L}\alpha_{l}(Y-g_{l}(X))^{2}-\sum_{l=1}^{L}\alpha_{l}(\hat{f}_{\text{ens}}(X)-g_{l}(X))^{2}.$ (39)
Proof.
$(Y-\hat{f}_{\text{ens}}(X))^{2}$ $=Y^{2}-2\hat{f}_{\text{ens}}(X)Y+\hat{f}_{\text{ens}}(X)^{2}$
$=\sum_{l=1}^{L}\alpha_{l}Y^{2}-2\sum_{l=1}^{L}\alpha_{l}g_{l}(X)Y+\hat{f}_{\text{ens}}(X)^{2}$
$=\sum_{l=1}^{L}\alpha_{l}Y^{2}-2\sum_{l=1}^{L}\alpha_{l}g_{l}(X)Y+\sum_{l=1}^{L}\alpha_{l}g_{l}(X)^{2}-\sum_{l=1}^{L}\alpha_{l}g_{l}(X)^{2}+\hat{f}_{\text{ens}}(X)^{2}$
$=\sum_{l=1}^{L}\alpha_{l}\left(Y^{2}-2g_{l}(X)Y+g_{l}(X)^{2}\right)-\left(\sum_{l=1}^{L}\alpha_{l}g_{l}(X)^{2}-\hat{f}_{\text{ens}}(X)^{2}\right)$
$=\sum_{l=1}^{L}\alpha_{l}\left(Y-g_{l}(X)\right)^{2}-\left(\sum_{l=1}^{L}\alpha_{l}g_{l}(X)^{2}-2\hat{f}_{\text{ens}}(X)^{2}+\hat{f}_{\text{ens}}(X)^{2}\right)$
$=\sum_{l=1}^{L}\alpha_{l}\left(Y-g_{l}(X)\right)^{2}-\left(\sum_{l=1}^{L}\alpha_{l}g_{l}(X)^{2}-2\sum_{l=1}^{L}\alpha_{l}g_{l}(X)\hat{f}_{\text{ens}}(X)+\hat{f}_{\text{ens}}(X)^{2}\right)$
$=\sum_{l=1}^{L}\alpha_{l}\left(Y-g_{l}(X)\right)^{2}-\sum_{l=1}^{L}\alpha_{l}\left(g_{l}(X)^{2}-2g_{l}(X)\hat{f}_{\text{ens}}(X)+\hat{f}_{\text{ens}}(X)^{2}\right)$
$=\sum_{l=1}^{L}\alpha_{l}\left(Y-g_{l}(X)\right)^{2}-\sum_{l=1}^{L}\alpha_{l}\left(\hat{f}_{\text{ens}}(X)-g_{l}(X)\right)^{2}$ (40)
B.5 Proof of Theorem 3
Now, we proof how subtraction in DeepBooTS mitigates concept drift, i.e., how Equation 9 diminishes the variance term in Equation 21. According to the estimation error $e_{l}\overset{i.i.d}{\sim}\mathcal{N}(0,\nu)$, we have $\text{Var}(g_{l}(X))=\nu$. Then, utilizing $\text{Cov}(g_{l},g_{k\neq l})=\mu$ in Theorem 1, we have the proof as follows:
Proof.
[…]
$=\frac{1}{\hbar^{2}}\text{Var}\left(\sum_{l=0}^{\hbar}\alpha g_{2l+1}(X)\right)+\frac{1}{\hbar^{2}}\text{Var}\left(\sum_{l=0}^{\hbar}\alpha g_{2l}(X)\right)-\frac{1}{\hbar^{2}}\text{Cov}\left(\sum_{l=0}^{\hbar}\alpha g_{2l+1}(X)\sum_{l=0}^{\hbar}\alpha g_{2l}(X)\right)$
$=\frac{1}{\hbar^{2}}\sum_{l=0}^{\hbar}\alpha^{2}\text{Var}\left(g_{2l+1}(X)\right)+\frac{1}{\hbar^{2}}\sum_{l=0}^{\hbar}\sum_{k=1,k\neq j}^{\hbar}\alpha^{2}\text{Cov}\left(g_{2l+1}(X)g_{2k+1}(X)\right)$
$\ \ \ \ \ \ +\frac{1}{\hbar^{2}}\sum_{l=0}^{\hbar}\alpha^{2}\text{Var}\left(g_{2l}(X)\right)+\frac{1}{\hbar^{2}}\sum_{l=0}^{\hbar}\sum_{k=1,k\neq j}^{\hbar}\alpha^{2}\text{Cov}\left(g_{2l}(X)g_{2k}(X)\right)$
$\ \ \ \ \ \ -\frac{1}{\hbar^{2}}\sum_{l=0}^{\hbar}\sum_{k=0}^{\hbar}\alpha^{2}\text{Cov}\left(g_{2l+1}(X)g_{2k}(X)\right)$
$=\frac{1}{\hbar^{2}}\sum_{l=0}^{\hbar}\alpha^{2}\nu+\frac{1}{\hbar^{2}}\sum_{l=0}^{\hbar}\sum_{k=1,k\neq j}^{\hbar}\alpha^{2}+\frac{1}{\hbar^{2}}\sum_{l=0}^{\hbar}\alpha^{2}\nu+\frac{1}{\hbar^{2}}\sum_{l=0}^{\hbar}\sum_{k=1,k\neq j}^{\hbar}\alpha^{2}\mu-\frac{1}{\hbar^{2}}\sum_{l=0}^{\hbar}\sum_{k=0}^{\hbar}\alpha^{2}\mu$
$=\frac{1}{\hbar}\alpha^{2}\nu+\frac{\hbar-1}{\hbar}\alpha^{2}\mu+\frac{1}{\hbar}\alpha^{2}\nu+\frac{\hbar-1}{\hbar}\alpha^{2}\mu-\alpha^{2}\mu$
$=\frac{2}{\hbar}\alpha^{2}\nu+2\frac{\hbar-1}{\hbar}\alpha^{2}\mu-\alpha^{2}\mu$
$<\frac{2}{\hbar}\alpha^{2}(\nu+\mu)$
$\leq\frac{4}{L}\alpha^{2}(\nu+\mu).$ (41)
[…]
Appendix C Dataset
C.1 Commonly used TS datasets
The information of the experiment datasets used in this paper are summarized as follows: (1) Electricity Transformer Temperature (ETT) dataset (Zhou2021Informer), which contains the data collected from two electricity transformers in two separated counties in China, including the load and the oil temperature recorded every 15 minutes (ETTm) or 1 hour (ETTh) between July 2016 and July 2018. (2) Electricity (ECL) dataset  [^1] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014 collects the hourly electricity consumption of 321 clients (each column) from 2012 to 2014. (3) Exchange (lai2018modeling) records the current exchange of 8 different countries from 1990 to 2016. (4) Traffic dataset  [^2] http://pems.dot.ca.gov records the occupation rate of freeway system across State of California measured by 861 sensors. (5) Weather dataset  [^3] https://www.bgc-jena.mpg.de/wetter records every 10 minutes for 21 meteorological indicators in Germany throughout 2020. (6) Solar-Energy (lai2018modeling) documents the solar power generation of 137 photovoltaic (PV) facilities in the year 2006, with data collected at 10-minute intervals. (7) The PEMS dataset (liu2022scinet) comprises publicly available traffic network data from California, collected within 5-minute intervals and encompassing 358 attributes. (8) Illness (ILI) dataset  [^4] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html describes the influenza-like illness patients in the United States between 2002 and 2021, which records the ratio of patients seen with illness and the total number of the patients. The detailed statistics information of the datasets is shown in Table 4.
Table 4: Details of the seven TS datasets.
Dataset length features frequency
ETTh1 17,420 7 1h
ETTh2 17,420 7 1h
ETTm1 69,680 7 15m
ETTm2 69,680 7 15m
Electricity 26,304 321 1h
Exchange 7,588 8 1d
Traffic 17,544 862 1h
Weather 52,696 21 10m
Solar 52,560 137 10m
PEMS 26,208 358 5m
Illness 966 7 7d
C.2 Monash TS Forecasting Datasets
(1) Saugeenday dataset  [^5] https://zenodo.org/records/4656058 contains a single very long time series representing the daily mean flow of the Saugeen River at Walkerton in cubic meters per second and the length of this time series is 23741. (2) Sunspot dataset  [^6] https://www.kaggle.com/datasets/robervalt/sunspots contains monthly numbers of sunspots, as from the World Data Center, aka SIDC, between 1749 and 2018 with a total observation of 3240 months. (3) M4 dataset (MAKRIDAKIS202054) is a collection of 100,000 time series used in the fourth Makridakis Prediction Contest. The dataset consists of a time series of annual, quarterly, monthly, and other weekly, daily, and hourly data. In this paper, we utilize the hourly version of the M4 dataset and standardize its length to 768. (4) NN5 dataset  [^7] http://www.neural-forecasting-competition.com/downloads/NN5/datasets/download.htm contains daily time series originated from the observation of daily withdrawals at 111 randomly selected different cash machines at different locations within England. (5) Oikolab Weather dataset  [^8] https://docs.oikolab.com contains hundreds of terabytes of weather data, all of them are post-processed data from national weather agencies. (6) US Births dataset  [^9] https://cran.r-project.org/web/packages/mosaicData provides birth rates and related data across the 50 states and DC from 2016 to 2021, with a total observation of 7305. (7) Rideshare dataset (godahewa2021monash) comprises diverse hourly time series representations of attributes pertinent to Uber and Lyft rideshare services across multiple locations in New York. The data spans from November 26th, 2018, to December 18th, 2018. All the above data sets are collected in the Monash library [^10] https://forecastingdata.org.
C.3 Large-Scale Time Series Datasets
Now we introduce three large-scale TS datasets:
(1) Milano (barlacchi2015multi) is provided by SKIL [^11] http://jol.telecomitalia.com/jolskil of Telecom Italia. The dataset was collected from November 1, 2013, to January 1, 2014, and the data is aggregated into 10-minute intervals over the whole city of Milan (62 days, 300 million records, about 19 GB). Milano records the time of the interaction and the specific radio base station that managed it. The dataset contains 10,000 nodes, and the length of each node is 1498. The ratio of training, validation and test datasets is 10:2:3.
(2) CBS is collected from all base station in a certain city in China, including up-link and down-link communication traffic data. The dataset is collected from June 8 to July 26, 2019, with a temporal interval of 60 minutes over the whole city (94 days, 51.9 million records, about 17.1 GB). CBS contains 4454 nodes, each with a length of 4032. The ratio of training, validation and test datasets is 3:1:1.
(3) C2TM (Modeling15Chen) makes an analysis on week-long traffic generated by a large population of people in a median-size city of China. C2TM analyses make use of request-response records extracted from traffic at the city scale, consisting of individuals’ activities during a continuous week (actually eight days from Aug. 19 to Aug. 26, 2012), with accurate timestamp and location information indicated by connected cellular base stations. C2TM contains 13269 nodes, each with a length of 192. The ratio of training, validation and test datasets is 5:1:2.
Appendix D Implementation Details
The model undergoes training utilizing the ADAM optimizer (kingma2014adam) and minimizing the Mean Squared Error (MSE) loss function. The training process is halted prematurely, typically within 10 epochs. The DeepBooTS architecture solely comprises the embedding layer and backbone architecture, devoid of any additional introduced hyperparameters. Refer to Appendix I for the hyperparameter sensitivity analysis. During model validation, two evaluation metrics are employed: MSE and Mean Absolute Error (MAE).
For the large-scale datasets, we initially employ the RSS algorithm in (liang2024act) to perform data sampling. In this algorithm, we random decompose the entire large graph into 17 and 25 sub-graphs for the CBS and Milano datasets, respectively. As in previous works (wang2024timexer; wang2024timemixer; liu2023itransformer; nie2022time; wu2021autoformer; Zhou2021Informer; liang2023does), the model is trained using the ADAM optimizer (kingma2014adam) under input-36-output-24 setting. The learning rate is set to 1e-4 with MSE loss. The batch size is set to 64 for training and 1 for testing, and the maximum number of epochs is set to 10.
Appendix E Performance Gains of DeepBooTS with Different Attention
Figure 9 breaks down how the choice of attention inside DeepBooTS affects forecast accuracy, reporting both MSE and MAE for five attention variants (Full-Attention, Prob-Attention, Autoformer, Flow-Attention, Period-Attention) on three TS datasets. On the highly stochastic Traffic series, vanilla dense attention still wins. Every sparse or structured variant increases error, with Periodic attention showing the largest drop‑off. On the Traffic dataset, ProbSparse and Periodic attention each cut MSE by about 3%, whereas Autoformer’s decomposition hurts performance heavily. On the Weather dataset, ProbSparse delivers the best MSE, Flow attention excels at MAE (–5.9%), and Autoformer again degrades accuracy. Periodic attention sits in the middle-better than Autoformer but slightly behind Full and ProbSparse in MSE.
Figure 9: Performance gains of DeepBooTS with different attention mechanisms.
Appendix F Performance Gains on Large-Scale TS Datasets
We also evaluated the performance gain when DeepBooTS is configured with other models, including PatchTST (nie2022time), STID (shao2022spatial), GWNet (wu2019graph), Autoformer (wu2021autoformer) and Informer (Zhou2021Informer). As shown in Fig. 10, DeepBooTS brings significant performance gains to other models in large-scale TS datasets. DeepBooTS consistently lowers forecasting error across all five backbones and three datasets, with relative gains ranging from small (2% on C2TM) to massive (up to 94% on CBS), highlighting its ability to stabilize and sharpen predictions even for models struggling under heavy distribution shifts.
Figure 10: Performance gains that DeepBooTS contributes to other models in large-scale TS datasets.
Appendix G Model Complexity and Computation Cost
[…]
Further, we evaluate the running time, memory usage, parameters, and FLOPs of our model in comparison to other Transformer-based models using identical settings. The computation costs are shown in the Table 5. It reveals that the proposed model has a smaller computational cost and faster running speed compared to others. It is noteworthy that the introduction of the auxiliary output branch only marginally increases the computational cost, i.e., the differences between the indicators of our model without gate and those of the original model are negligible.
Table 5: Model complexity, training efficiency and computation cost.
Models Seconds/Epoch GPU Memory Usage (GB) Parameters (MB) FLOPs (GB)
DeepBooTS 28.7 1.2 8.8 3.1
DeepBooTS w/o Gate 28.0 1.1 6.5 2.5
PatchTST 30.1 2.3 8.4 17.4
FEDformer 396 7.1 14.7 558.1
Autoformer 84.5 4.8 10.5 66.2
Informer 81.8 2.3 11.3 63.3
Appendix H Concept Drift in TS Forecasting
We recorded the training process and verification errors of various models, including Informer, Autoformer, FEDformer, DLinear, iTransformer, and DeepBooTS (ours). The experimental setup is the same as in the original paper of these methods. As shown in Fig. 11 (zoom in for better viewing), the model exhibits a significant concept drift phenomenon, particularly in the early stages of training. We analyze that the main cause of concept drift in this scenario is the high model complexity: Using highly complex models with many parameters can lead to the model “memorizing” the noise and specific details in the training data. This results in excellent performance on the training data but poor generalization to new data. In this case, using a linear model (e.g., DLinear in Fig. 11(e)) will mitigate concept drift but will also fail to capture temporal correlations. The proposed DeepBooTS method adopts the idea of ensemble learning to provide flexible complexity for the model, thereby alleviating the poor generalization esulting from concept drift the model to the training data. This can be verified in Fig. 11(f), where DeepBooTS has high training error but low test error.
Figure 11: Concept drift in TS Forecasting. Experiments are conducted utilizing various models on the Traffic dataset. Similar phenomena can be observed on other datasets.
Appendix I Hyperparameter Sensitivity
We evaluate the hyperparameter sensitivity of DeepBooTS with respect to the learning rate, the number of the block, the batch size and the embedding dimension. As shown in Fig. 12, the performance of DeepBooTS fluctuates under different hyperparameter settings. In most cases, increasing the number of blocks tends to enhance model performance. Once again, this confirms that DeepBooTS exhibits resilience against concept drift across diverse datasets. Notably, we observe that the learning rate, being the most prevalent influencing factor, especially in scenarios involving numerous attributes. Meanwhile, modifying the batch size induces minor fluctuations in the model’s performance, albeit with a limited impact. Furthermore, we conducted extensive analysis on hyperparameters, as shown in Tables 6, 7, and 8. It is evident that the proposed model is less insensitive to numerous hyperparameters. In summary, DeepBooTS exhibits low sensitivity to these hyperparameters, thereby enhancing its resilience against concept drift.
Figure 12: Hyperparameter sensitivity with respect to the number of block, the learning rate and the number of batch size. The results are recorded with the input length $I=96$ and the prediction length $O=12$ for PEMS and $O=96$ for others.
Table 6: Ablation studies of DeepBooTS’s hyperparameters on the Electricity dataset.
Block Number Embedding Dimension Learning Rate
Metrics 2 3 4 5 6 92 384 512 768 1E-3 5E-4 1E-4 5E-5
MSE 0.137 0.137 0.137 0.138 0.137 0.140 0.137 0.143 0.136 0.143 0.143 0.144 0.145
MAE 0.231 0.231 0.231 0.230 0.231 0.233 0.231 0.235 0.231 0.231 0.231 0.230 0.232
RMSP 0.214 0.214 0.256 0.215 0.213 0.219 0.215 0.214 0.212 0.214 0.214 0.212 0.226
MAPE 2.295 2.294 2.34 2.32 2.32 2.397 2.318 2.295 2.305 2.29 2.303 2.302 2.35
sMAPE 0.466 0.464 0.467 0.466 0.464 0.475 0.467 0.466 0.468 0.466 0.466 0.465 0.469
MASE 0.257 0.256 0.259 0.257 0.257 0.266 0.258 0.257 0.261 0.257 0.257 0.258 0.275
Q25 0.232 0.229 0.232 0.232 0.231 0.234 0.232 0.232 0.233 0.232 0.229 0.229 0.238
Q75 0.23 0.232 0.23 0.231 0.23 0.24 0.231 0.23 0.231 0.23 0.232 0.233 0.243
Table 7: Ablation studies of DeepBooTS’s hyperparameters on the Traffic dataset.
Block Number Embedding Dimension Learning Rate
Metrics 2 3 4 5 6 92 384 512 768 1E-3 5E-4 1E-4 5E-5
MSE 0.400 0.401 0.400 0.400 0.401 0.393 0.387 0.387 0.389 0.389 0.391 0.399 0.404
MAE 0.266 0.267 0.267 0.267 0.267 0.265 0.258 0.259 0.257 0.257 0.256 0.261 0.271
RMSP 0.208 0.209 0.208 0.209 0.208 0.207 0.198 0.199 0.200 0.207 0.203 0.208 0.214
MAPE 2.730 2.770 2.765 2.750 2.750 2.760 2.670 2.690 2.700 2.720 2.660 2.720 2.820
sMAPE 0.485 0.486 0.486 0.486 0.486 0.481 0.470 0.480 0.480 0.484 0.482 0.491 0.492
MASE 0.257 0.258 0.258 0.258 0.258 0.255 0.248 0.250 0.252 0.249 0.250 0.274 0.262
Q25 0.262 0.265 0.264 0.264 0.264 0.263 0.256 0.256 0.254 0.253 0.255 0.259 0.268
Q75 0.271 0.270 0.270 0.270 0.271 0.267 0.260 0.262 0.261 0.261 0.258 0.264 0.275
Table 8: Ablation studies of DeepBooTS’s hyperparameters on the Weather dataset.
Block Number Embedding Dimension Learning Rate
Metrics 2 3 4 5 6 92 384 512 768 1E-3 5E-4 1E-4 5E-5
MSE 0.160 0.156 0.160 0.159 0.157 0.161 0.156 0.158 0.159 0.215 0.168 0.157 0.159
MAE 0.203 0.201 0.206 0.203 0.199 0.205 0.201 0.202 0.202 0.271 0.211 0.202 0.202
RMSP 0.213 0.213 0.220 0.215 0.206 0.213 0.211 0.210 0.212 0.324 0.218 0.212 0.210
MAPE 8.990 8.860 9.220 9.430 9.130 10.600 9.151 8.387 8.450 17.680 8.530 7.900 7.860
sMAPE 0.529 0.524 0.534 0.528 0.519 0.531 0.526 0.525 0.524 0.637 0.542 0.529 0.523
MASE 1.162 1.110 1.172 1.161 1.090 1.263 1.171 1.151 1.094 1.230 1.202 1.170 1.200
Q25 0.198 0.196 0.203 0.197 0.188 0.197 0.195 0.197 0.189 0.272 0.206 0.200 0.195
Q75 0.209 0.207 0.209 0.209 0.210 0.212 0.208 0.206 0.215 0.269 0.215 0.205 0.209
Appendix J Different Attention Promotion
Figure 13: Visualization of Attention utilized in DeepBooTS. The model is trained on Traffic dataset with 862 attributes under the setting of Input-96-Predict-96. Both the Attention score (top) and the post-softmax score (below) of the 8 heads (H) are from the first block.
We visualized the Attention maps for all Attention heads in the initial layer of DeepBooTS, and the results are presented in Fig. 13. It is evident that the Attention score graph exhibits numerous bar structures, which is especially prominent on the post-softmax Attention map. This implies that there exists a row in Query that bears a striking resemblance to all the columns in Key. This scenario arises when there are numerous attributes, many of which are homogeneously represented. Our speculation suggests that this capability of the Attention may explain its ability to capture subtle patterns in TS without succumbing to concept drift. Furthermore, we substituted its original Attention with other novel Attention mechanisms to compare the resulting changes in model performance. The full results for all prediction lengths are provided in Table 9.
Table 9: Ablation of different Attention layers.
Attention Layer FullAttention ProbAttention AutoCorrelation FlowAttention PeriodAttention
Dataset Length MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Traffic 96 0.386 0.258 0.390 0.259 0.396 0.261 0.385 0.256 0.414 0.273
192 0.398 0.263 0.410 0.268 0.413 0.268 0.406 0.265 0.431 0.279
336 0.409 0.270 0.425 0.274 0.429 0.275 0.424 0.274 0.447 0.286
720 0.431 0.287 0.459 0.293 0.461 0.293 0.459 0.293 0.478 0.303
Avg 0.406 0.270 0.421 0.274 0.425 0.274 0.418 0.272 0.443 0.285
Electricity 96 0.135 0.229 0.133 0.226 0.174 0.256 0.134 0.228 0.142 0.233
192 0.151 0.246 0.148 0.241 0.179 0.263 0.150 0.251 0.148 0.243
336 0.172 0.268 0.167 0.263 0.195 0.279 0.174 0.267 0.166 0.262
720 0.204 0.294 0.195 0.290 0.232 0.311 0.193 0.294 0.194 0.289
Avg 0.166 0.259 0.161 0.255 0.195 0.277 0.163 0.260 0.162 0.258
Weather 96 0.160 0.206 0.159 0.204 0.179 0.220 0.169 0.209 0.170 0.212
192 0.208 0.250 0.207 0.248 0.228 0.261 0.220 0.255 0.214 0.253
336 0.269 0.296 0.265 0.291 0.288 0.304 0.276 0.294 0.273 0.296
720 0.343 0.334 0.340 0.342 0.367 0.356 0.354 0.348 0.352 0.347
Avg 0.245 0.271 0.243 0.271 0.266 0.285 0.255 0.276 0.252 0.277
    The input length $I$ is set as 96, while the prediction lengths $O\in$ {96, 192, 336, 720}.
Appendix K Full Results on ETT Datasets
The ETT dataset records electricity data of four different granularities and types. We offer an in-depth comparison of DeepBooTS utilizing the complete ETT dataset to facilitate future research endeavors. Detailed results are provided in Table 10. It is evident that DeepBooTS demonstrates excellent performance on the complete ETT dataset.
Table 10: Full Multivariate Forecasting Results on ETT dataset.
Models DeepBooTS Periodformer FEDformer Autoformer Informer LogTrans Reformer
Length MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.370 0.394 0.375 0.395 0.395 0.424 0.449 0.459 0.865 0.713 0.878 0.74 0.837 0.728
192 0.423 0.427 0.413 0.421 0.469 0.47 0.5 0.482 1.008 0.792 1.037 0.824 0.923 0.766
336 0.465 0.446 0.443 0.441 0.530 0.499 0.521 0.496 1.107 0.809 1.238 0.932 1.097 0.835
720 0.465 0.464 0.467 0.469 0.598 0.544 0.514 0.512 1.181 0.865 1.135 0.852 1.257 0.889
Avg 0.431 0.433 0.425 0.432 0.498 0.484 0.496 0.487 1.040 0.795 1.072 0.837 1.029 0.805
ETTh2 96 0.285 0.333 0.313 0.356 0.394 0.414 0.358 0.397 3.755 1.525 2.116 1.197 2.626 1.317
192 0.354 0.377 0.389 0.405 0.439 0.445 0.456 0.452 5.602 1.931 4.315 1.635 11.12 2.979
336 0.405 0.420 0.418 0.432 0.482 0.48 0.482 0.486 4.721 1.835 1.124 1.604 9.323 2.769
720 0.418 0.438 0.427 0.444 0.5 0.509 0.515 0.511 3.647 1.625 3.188 1.54 3.874 1.697
Avg 0.366 0.392 0.387 0.409 0.454 0.462 0.453 0.462 4.431 1.729 2.686 1.494 6.736 2.191
ETTm1 96 0.311 0.340 0.337 0.378 0.378 0.418 0.505 0.475 0.672 0.571 0.600 0.546 0.538 0.528
192 0.363 0.379 0.413 0.431 0.464 0.463 0.553 0.496 0.795 0.669 0.837 0.7 0.658 0.592
336 0.393 0.391 0.428 0.441 0.508 0.487 0.621 0.537 1.212 0.871 1.124 0.832 0.898 0.721
720 0.454 0.442 0.483 0.483 0.561 0.515 0.671 0.561 1.166 0.823 1.153 0.82 1.102 0.841
Avg 0.380 0.388 0.415 0.433 0.478 0.471 0.588 0.517 0.961 0.734 0.929 0.725 0.799 0.671
ETTm2 96 0.171 0.250 0.186 0.274 0.204 0.288 0.255 0.339 0.365 0.453 0.768 0.642 0.658 0.619
192 0.235 0.292 0.252 0.317 0.316 0.363 0.281 0.34 0.533 0.563 0.989 0.757 1.078 0.827
336 0.293 0.330 0.311 0.355 0.359 0.387 0.339 0.372 1.363 0.887 1.334 0.872 1.549 0.972
720 0.393 0.389 0.402 0.405 0.433 0.432 0.422 0.419 3.379 1.338 3.048 1.328 2.631 1.242
Avg 0.273 0.315 0.288 0.338 0.328 0.368 0.324 0.368 1.410 0.810 1.535 0.900 1.479 0.915
$1^{\text{st}}$ Count 17 17 3 3 0 0 0 0 0 0 0 0 0 0
Appendix L Pseudocode of DeepBooTS
To facilitate a comprehensive understanding of DeepBooTS’s working principle, we offer detailed pseudocode outlining its implementation, as shown in Algorithm 1. The implementation presented here outlines the core ideas of DeepBooTS. It is evident that the deployment procedure of DeepBooTS exhibits a relative simplicity, characterized by the inclusion of several iteratively applied blocks. This property renders it highly versatile for integrating newly devised Attention mechanisms or modules. As demonstrated in Appendix J, the substitution of Attention mechanisms in DeepBooTS with novel alternatives yields superior generalization.
Algorithm 1 DeepBooTS Architecture.
0: Batch size $B$, input lookback time series $\mathbf{X}\in\mathbb{R}^{B\times I\times D}$; input length II; predicted length $O$; embedding dimension $E$; the number of the block $L$; the output length $H$ in each block.
[…]
22:End for
[…]
25:for $\mathbf{O}_{L}=\texttt{Linear}(\mathbf{O}_{L-1})$ $\mathbf{O}_{L}\in\mathbb{R}^{B\times D\times O}$}
[…]
Appendix M Full Multivariate TS Forecasting Results
The full results for multivariate TS forecasting are presented in Table 11. It is evident that DeepBooTS demonstrates excellent performance on the complete multivariate TS forecasting. Notably, several pioneering models have also achieved competitive performance on certain datasets under particular settings. For instance, Informer, considered a groundbreaking model in long-term TS forecasting, demonstrates advanced performance on the Solar-Energy dataset with input-96-predict-192 and -720 settings. This is due to the substantial presence of zero values on each column attribute of the Solor-Energy dataset. This renders the KL-divergence based ProbSparse Attention, as adopted in Informer, highly effective on this sparse dataset. Additionally, linear-based methods (e.g., DLinear) have demonstrated promising results on the Weather dataset with input-336-predict-720 setting, while convolution-based methods (e.g., SCINet) have yielded favorable results on the PEMS dataset with input-168-predict-192 setting. This phenomenon can be ascribed to a twofold interplay of factors. Previously, the diversity of input settings exerts a direct influence on model generalization. Secondarily, other models exhibit a propensity to overfit non-stationary TS characterized by aperiodic fluctuations. Remarkably, DeepBooTS adeptly mitigates both concept drift and underfitting challenges in multivariate TS forecasting, thereby enhancing its overall performance. Particularly on datasets with numerous attributes, e.g., Traffic and Solor-Energy, DeepBooTS achieves superior performance by feeding the learned meaningful patterns to the output layer at each block.
Table 11: Multivariate TS forecasting results on six benchmark datasets.
Model DeepBooTS* DeepBooTS TimeXer TimeMixer iTransformer Crossformer PatchTST DLinear FEDformer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETT 96 0.280 0.336 0.284 0.329 0.280 0.338 0.290 0.339 0.297 0.349 0.745 0.584 0.299 0.347 0.333 0.387 0.358 0.397
192 0.332 0.371 0.344 0.369 0.348 0.377 0.350 0.373 0.380 0.400 0.877 0.656 0.365 0.404 0.477 0.476 0.429 0.439
336 0.362 0.392 0.389 0.397 0.393 0.404 0.390 0.404 0.428 0.432 1.043 0.731 0.413 0.415 0.594 0.541 0.496 0.487
720 0.410 0.429 0.433 0.433 0.430 0.432 0.439 0.438 0.427 0.445 1.104 0.763 0.458 0.451 0.831 0.657 0.463 0.474
Avg 0.346 0.382 0.362 0.382 0.365 0.388 0.367 0.389 0.383 0.407 0.942 0.684 0.384 0.404 0.559 0.515 0.437 0.449
Traffic 96 0.350 0.250 0.386 0.258 0.428 0.271 0.462 0.285 0.395 0.268 0.522 0.290 0.526 0.347 0.650 0.396 0.587 0.366
192 0.375 0.261 0.398 0.263 0.448 0.282 0.473 0.296 0.417 0.276 0.530 0.293 0.522 0.332 0.598 0.370 0.604 0.373
336 0.381 0.264 0.409 0.270 0.473 0.289 0.498 0.296 0.433 0.283 0.558 0.305 0.517 0.334 0.605 0.373 0.621 0.383
720 0.386 0.268 0.431 0.287 0.516 0.307 0.506 0.313 0.467 0.302 0.589 0.328 0.552 0.352 0.645 0.394 0.626 0.382
Avg 0.373 0.261 0.406 0.270 0.466 0.287 0.484 0.297 0.428 0.282 0.550 0.304 0.529 0.341 0.625 0.383 0.610 0.376
ELC 96 0.128 0.223 0.135 0.229 0.140 0.242 0.153 0.247 0.148 0.240 0.219 0.314 0.190 0.296 0.197 0.282 0.193 0.308
192 0.148 0.24 0.151 0.246 0.157 0.256 0.166 0.256 0.162 0.253 0.231 0.322 0.199 0.304 0.196 0.285 0.201 0.315
336 0.164 0.26 0.172 0.268 0.176 0.275 0.185 0.277 0.178 0.269 0.246 0.337 0.217 0.319 0.209 0.301 0.214 0.329
720 0.192 0.284 0.204 0.294 0.211 0.306 0.225 0.310 0.225 0.317 0.280 0.363 0.258 0.352 0.245 0.333 0.246 0.355
Avg 0.158 0.252 0.166 0.259 0.171 0.270 0.182 0.272 0.178 0.270 0.244 0.334 0.216 0.318 0.212 0.300 0.214 0.327
Weather 96 0.150 0.201 0.160 0.206 0.157 0.205 0.163 0.209 0.174 0.214 0.158 0.230 0.186 0.227 0.196 0.255 0.217 0.296
192 0.194 0.244 0.208 0.250 0.204 0.247 0.208 0.250 0.221 0.254 0.206 0.277 0.234 0.265 0.237 0.296 0.276 0.336
336 0.245 0.282 0.269 0.296 0.261 0.290 0.251 0.287 0.278 0.296 0.272 0.335 0.284 0.301 0.283 0.335 0.339 0.380
720 0.320 0.336 0.343 0.334 0.340 0.341 0.339 0.341 0.358 0.349 0.398 0.418 0.356 0.349 0.345 0.381 0.403 0.428
Avg 0.227 0.266 0.245 0.271 0.241 0.271 0.240 0.271 0.258 0.279 0.259 0.315 0.265 0.285 0.265 0.317 0.309 0.360
Solar 96 0.181 0.222 0.192 0.222 0.229 0.273 0.189 0.259 0.203 0.237 0.310 0.331 0.265 0.323 0.290 0.378 0.242 0.342
192 0.198 0.239 0.230 0.251 0.265 0.297 0.222 0.283 0.233 0.261 0.734 0.725 0.288 0.332 0.320 0.398 0.285 0.380
336 0.202 0.245 0.243 0.263 0.293 0.312 0.231 0.292 0.248 0.273 0.750 0.735 0.301 0.339 0.353 0.415 0.282 0.376
720 0.206 0.252 0.243 0.265 0.302 0.318 0.223 0.285 0.249 0.275 0.769 0.765 0.295 0.336 0.356 0.413 0.357 0.427
Avg 0.197 0.240 0.227 0.250 0.272 0.300 0.216 0.280 0.233 0.262 0.641 0.639 0.287 0.333 0.330 0.401 0.291 0.381
PEMS 12 0.057 0.157 0.067 0.171 0.092 0.193 0.078 0.187 0.071 0.174 0.090 0.203 0.099 0.216 0.122 0.243 0.126 0.251
24 0.070 0.173 0.093 0.203 0.111 0.218 0.108 0.216 0.093 0.201 0.121 0.240 0.142 0.259 0.201 0.317 0.149 0.275
36 0.083 0.186 0.125 0.237 0.135 0.257 0.151 0.254 0.125 0.236 0.202 0.317 0.211 0.319 0.333 0.425 0.227 0.348
48 0.091 0.195 0.151 0.262 0.226 0.232 0.213 0.311 0.160 0.270 0.262 0.367 0.269 0.370 0.457 0.515 0.348 0.434
Avg 0.075 0.178 0.109 0.218 0.141 0.225 0.138 0.242 0.113 0.221 0.169 0.281 0.180 0.291 0.278 0.375 0.213 0.327
$1^{st}$ or $2^{st}$ 30 30 15 22 3 4 8 2 3 2 3 0 0 0 1 0 0 0
Appendix N Full Univariate TS Forecasting Results
The full results for univariate TS forecasting are presented in Table 12. As other models, e.g., iTransformer (liu2023itransformer), PatchTST (nie2022time) and Crossformer (zhang2022crossformer) do not offer performance information for all prediction lengths, we compare our method with those that provide comprehensive performance analysis, including Periodformer (liang2023does), FEDformer (zhou2022fedformer), Autoformer (wu2021autoformer), Informer (Zhou2021Informer), LogTrans (li2019enhancing) and Reformer (Kitaev2020Reformer). Despite Periodformer being a model that determines optimal hyperparameters through search, the proposed method outperforms benchmark approaches by achieving the highest count of leading terms across various prediction lengths. This reaffirms the effectiveness of DeepBooTS.
Table 12: Univariate TS forecasting results on benchmark datasets.
Model DeepBooTS Periodformer FEDformer Autoformer Informer LogTrans Reformer
Length MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.055 0.177 0.068 0.203 0.079 0.215 0.071 0.206 0.193 0.377 0.283 0.468 0.532 0.569
192 0.072 0.204 0.088 0.228 0.104 0.245 0.114 0.262 0.217 0.395 0.234 0.409 0.568 0.575
336 0.08 0.219 0.105 0.256 0.119 0.270 0.107 0.258 0.202 0.381 0.386 0.546 0.635 0.589
720 0.079 0.224 0.109 0.262 0.142 0.299 0.126 0.283 0.183 0.355 0.475 0.628 0.762 0.666
Avg 0.072 0.206 0.093 0.237 0.111 0.257 0.105 0.252 0.199 0.377 0.345 0.513 0.624 0.600
ETTh2 96 0.129 0.275 0.125 0.272 0.128 0.271 0.153 0.306 0.213 0.373 0.217 0.379 1.411 0.838
192 0.178 0.329 0.175 0.329 0.185 0.33 0.204 0.351 0.227 0.387 0.281 0.429 5.658 1.671
336 0.211 0.365 0.219 0.372 0.231 0.378 0.246 0.389 0.242 0.401 0.293 0.437 4.777 1.582
720 0.220 0.377 0.249 0.400 0.278 0.42 0.268 0.409 0.291 0.439 0.218 0.387 2.042 1.039
Avg 0.185 0.337 0.192 0.343 0.206 0.350 0.218 0.364 0.243 0.400 0.252 0.408 3.472 1.283
ETTm1 96 0.029 0.126 0.033 0.139 0.033 0.140 0.056 0.183 0.109 0.277 0.049 0.171 0.296 0.355
192 0.044 0.158 0.052 0.177 0.058 0.186 0.081 0.216 0.151 0.310 0.157 0.317 0.429 0.474
336 0.057 0.185 0.070 0.267 0.084 0.231 0.076 0.218 0.427 0.591 0.289 0.459 0.585 0.583
720 0.080 0.218 0.081 0.221 0.102 0.250 0.110 0.267 0.438 0.586 0.430 0.579 0.782 0.73
Avg 0.052 0.172 0.059 0.201 0.069 0.202 0.081 0.221 0.281 0.441 0.231 0.382 0.523 0.536
ETTm2 96 0.064 0.180 0.060 0.182 0.063 0.189 0.065 0.189 0.08 0.217 0.075 0.208 0.077 0.214
192 0.099 0.233 0.099 0.236 0.110 0.252 0.118 0.256 0.112 0.259 0.129 0.275 0.138 0.290
336 0.129 0.273 0.129 0.275 0.147 0.301 0.154 0.305 0.166 0.314 0.154 0.302 0.160 0.313
720 0.180 0.329 0.170 0.317 0.219 0.368 0.182 0.335 0.228 0.380 0.160 0.322 0.168 0.334
Avg 0.118 0.254 0.115 0.253 0.135 0.278 0.130 0.271 0.147 0.293 0.130 0.277 0.136 0.288
Traffic 96 0.127 0.202 0.143 0.222 0.170 0.263 0.246 0.346 0.257 0.353 0.226 0.317 0.313 0.383
192 0.135 0.211 0.146 0.227 0.173 0.265 0.266 0.37 0.299 0.376 0.314 0.408 0.386 0.453
336 0.130 0.215 0.147 0.231 0.178 0.266 0.263 0.371 0.312 0.387 0.387 0.453 0.423 0.468
720 0.135 0.218 0.164 0.252 0.187 0.286 0.269 0.372 0.366 0.436 0.437 0.491 0.378 0.433
Avg 0.132 0.212 0.150 0.233 0.177 0.270 0.261 0.365 0.309 0.388 0.341 0.417 0.375 0.434
ELC 96 0.249 0.358 0.236 0.349 0.262 0.378 0.341 0.438 0.258 0.367 0.288 0.393 0.275 0.379
192 0.286 0.379 0.277 0.369 0.316 0.410 0.345 0.428 0.285 0.388 0.432 0.483 0.304 0.402
336 0.337 0.413 0.324 0.400 0.361 0.445 0.406 0.470 0.336 0.423 0.430 0.483 0.37 0.448
720 0.385 0.454 0.353 0.437 0.448 0.501 0.565 0.581 0.607 0.599 0.491 0.531 0.46 0.511
Avg 0.314 0.401 0.298 0.389 0.347 0.434 0.414 0.479 0.372 0.444 0.410 0.473 0.352 0.435
Weather 96 0.0012 0.0263 0.0015 0.0300 0.0035 0.046 0.0110 0.081 0.004 0.044 0.0046 0.052 0.012 0.087
192 0.0014 0.0283 0.0015 0.0307 0.0054 0.059 0.0075 0.067 0.002 0.040 0.006 0.060 0.0098 0.044
336 0.0015 0.0294 0.0017 0.0313 0.008 0.072 0.0063 0.062 0.004 0.049 0.006 0.054 0.013 0.100
720 0.002 0.0333 0.0020 0.0348 0.015 0.091 0.0085 0.070 0.003 0.042 0.007 0.059 0.011 0.083
Avg 0.0015 0.0293 0.0017 0.0317 0.008 0.067 0.0083 0.0700 0.0033 0.0438 0.0059 0.0563 0.0115 0.0785
Exchange 96 0.096 0.226 0.092 0.226 0.131 0.284 0.241 0.387 1.327 0.944 0.237 0.377 0.298 0.444
192 0.200 0.332 0.198 0.341 0.277 0.420 0.300 0.369 1.258 0.924 0.738 0.619 0.777 0.719
336 0.400 0.473 0.370 0.471 0.426 0.511 0.509 0.524 2.179 1.296 2.018 1.0700 1.833 1.128
720 1.020 0.779 0.753 0.696 1.162 0.832 1.260 0.867 1.28 0.953 2.405 1.175 1.203 0.956
Avg 0.429 0.453 0.353 0.434 0.499 0.512 0.578 0.537 1.511 1.029 1.350 0.810 1.028 0.812
$1^{\text{st}}$ Count 24 27 14 12 0 1 0 0 0 0 2 0 0 0
    $\diamondsuit$} denotes the maximum search range of the input length.
Appendix O Comparative Analysis
Notably, several pioneering models have also achieved competitive performance on certain datasets under particular settings. For instance, Informer, considered a groundbreaking model in long-term TS forecasting, demonstrates advanced performance on the Solar-Energy dataset with input-96-predict-192 and -720 settings. This is due to the substantial presence of zero values on each column attribute of the Solor-Energy dataset. This renders the KL-divergence based ProbSparse Attention, as adopted in Informer, highly effective on this sparse dataset. Additionally, linear-based methods (e.g., DLinear) have demonstrated promising results on the Weather dataset with input-336-predict-720 setting, while convolution-based methods (e.g., SCINet) have yielded favorable results on the PEMS dataset with input-168-predict-192 setting. This phenomenon can be ascribed to a twofold interplay of factors. Previously, the diversity of input settings exerts a direct influence on model generalization. Secondarily, other models exhibit a propensity to overfit non-stationary TS characterized by aperiodic fluctuations. Remarkably, DeepBooTS adeptly mitigates both concept drift and underfitting challenges in multivariate TS forecasting, thereby enhancing its overall performance. Particularly on datasets with numerous attributes, e.g., Traffic and Solor-Energy, DeepBooTS achieves superior performance by feeding the learned meaningful patterns to the output layer at each block.
Appendix P Interpretability of DeepBooTS
Fig. 14 illustrates that DeepBooTS learns a fully interpretable, hierarchical decomposition of the target series by assigning distinct temporal patterns to each block. In the three‐block, low‐dimensional setting, Block 1 captures the dominant seasonal cycle, Block 2 models the residual phase and amplitude errors, and Block 3 provides a constant bias correction; increasing the embedding dimension sharpens the periodic waveforms without altering this structure. When extended to five blocks at high dimensionality, Blocks 1 and 2 isolate the two strongest periodicities, Blocks 3 and 4 extract mid‐frequency fluctuations, and Block 5 again serves as a global baseline. Summing these block outputs reproduces the final forecast almost exactly, demonstrating DeepBooTS’s ability to decompose complex series into interpretable, scale‐specific components and thereby facilitate both efficient learning and transparent model analysis.
Figure 14: The visualization of each block in DeepBooTS.
Appendix Q Evaluation Metrics
This paper uses a variety of metrics, including MSE (Mean Square Error), MAE (Mean Absolute Error), RMSP (Root Median Square Percent), MAPE (Mean Absolute Percentage Error), sMAPE (symmetric MAPE), MASE (Mean Absolute Scaled Error) and Quantile loss. All evaluation metrics adopted in this paper are as follows:
[…]
$\text{QuantileLoss}(Y,\hat{Y},q)$ $=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}_{\hat{y}_{i}\geq y_{i}}(1-q)|y_{i}-\hat{y}_{i}|+\mathbb{I}_{\hat{y}_{i}<y_{i}}q|y_{i}-\hat{y}_{i}|.$ (48)
Appendix R Full TS forecasting on Monash’s Repository
The full results for mulvariate and univariate TS forecasting on 4 commonly used datasets and 7 Monash TS datasets are presented in Table 13. Due to different characteristics of the dataset, the output lengths Y1, Y2, Y3 and Y4, are slightly different. Specifically, for Traffic, Electricity, Solar, OiK_Weather, Sunspot, Saugeenday and M4, the input length is 96, and the output lengths are 96, 192, 336, and 720, respectively. For ILI dataset, the input length is 36, and the output lengths are 24, 36, 48, and 60, respectively. For NN5 and Rideshare datasets, the input length is 48, and the output lengths are 12, 24, 36, and 48, respectively. It is evident that the proposed method outperforms benchmark approaches by achieving the highest count of leading terms across various prediction lengths. This reaffirms the effectiveness of DeepBooTS.
Table 13: Mulvariate and univariate forecasting results with diverse metrics on Monash TS datasets.
Mulvariate DeepBooTS iTransformer DLinear Autoformer Informer
Dataset Metric Y1 Y2 Y3 Y4 Avg IMP Y1 Y2 Y3 Y4 Avg Y1 Y2 Y3 Y4 Avg Y1 Y2 Y3 Y4 Avg Y1 Y2 Y3 Y4 Avg
Traffic MSE 0.386 0.398 0.409 0.431 0.406 5.14% 0.395 0.417 0.433 0.467 0.428 0.650 0.598 0.605 0.645 0.625 0.613 0.616 0.622 0.660 0.628 0.719 0.696 0.777 0.864 0.764
MAE 0.258 0.263 0.270 0.287 0.270 4.26% 0.268 0.276 0.283 0.302 0.282 0.396 0.370 0.373 0.394 0.383 0.388 0.382 0.337 0.408 0.379 0.391 0.379 0.420 0.472 0.416
RMSP 0.205 0.200 0.207 0.231 0.211 6.64% 0.214 0.219 0.225 0.247 0.226 0.350 0.317 0.322 0.356 0.336 0.437 0.409 0.384 0.432 0.416 0.423 0.416 0.522 0.715 0.519
MAPE 2.657 2.694 2.715 2.797 2.716 8.24% 2.885 2.955 2.942 3.059 2.960 5.018 4.477 4.290 4.196 4.495 4.377 4.198 4.254 4.288 4.279 4.760 4.677 6.115 7.016 5.642
sMAPE 0.479 0.476 0.485 0.514 0.489 2.59% 0.485 0.494 0.501 0.529 0.502 0.662 0.625 0.629 0.663 0.645 0.721 0.687 0.657 0.718 0.696 0.718 0.712 0.810 0.978 0.805
MASE 0.256 0.257 0.264 0.285 0.266 1.12% 0.255 0.263 0.268 0.288 0.269 0.397 0.365 0.366 0.390 0.380 0.405 0.390 0.364 0.401 0.390 0.415 0.406 0.468 0.595 0.471
Q25 0.255 0.264 0.268 0.281 0.267 4.64% 0.271 0.279 0.279 0.290 0.280 0.380 0.357 0.357 0.373 0.367 0.440 0.414 0.395 0.419 0.417 0.397 0.372 0.461 0.614 0.461
Q75 0.270 0.270 0.280 0.304 0.281 1.40% 0.267 0.276 0.286 0.311 0.285 0.412 0.383 0.390 0.417 0.401 0.391 0.387 0.374 0.416 0.392 0.456 0.465 0.536 0.621 0.520
Electricity MSE 0.143 0.162 0.179 0.204 0.172 3.37% 0.148 0.162 0.178 0.225 0.178 0.197 0.196 0.209 0.245 0.212 0.201 0.222 0.231 0.254 0.227 0.274 0.296 0.300 0.373 0.311
MAE 0.235 0.253 0.271 0.294 0.263 2.59% 0.240 0.253 0.269 0.317 0.270 0.282 0.285 0.301 0.333 0.300 0.317 0.334 0.338 0.361 0.338 0.368 0.386 0.394 0.439 0.397
RMSP 0.215 0.230 0.253 0.287 0.246 2.38% 0.223 0.239 0.256 0.290 0.252 0.310 0.316 0.334 0.370 0.333 0.324 0.346 0.367 0.374 0.353 0.429 0.796 0.487 0.680 0.598
MAPE 2.296 2.431 2.699 3.194 2.655 2.10% 2.493 2.732 2.713 2.910 2.712 2.695 2.715 2.667 2.760 2.709 3.515 3.309 3.626 3.401 3.463 3.789 3.354 2.354 3.826 3.331
sMAPE 0.466 0.487 0.514 0.558 0.506 1.75% 0.479 0.498 0.519 0.564 0.515 0.598 0.601 0.621 0.660 0.620 0.584 0.619 0.639 0.642 0.621 0.711 1.148 0.786 0.988 0.908
MASE 0.257 0.276 0.293 0.321 0.287 3.37% 0.267 0.284 0.298 0.340 0.297 0.357 0.359 0.374 0.405 0.374 0.329 0.359 0.380 0.379 0.362 0.414 0.832 0.445 0.614 0.576
Q25 0.232 0.239 0.270 0.321 0.266 -1.53% 0.232 0.255 0.269 0.290 0.262 0.314 0.317 0.327 0.356 0.329 0.349 0.342 0.378 0.384 0.363 0.406 0.665 0.421 0.590 0.521
Q75 0.230 0.254 0.261 0.276 0.255 7.27% 0.250 0.258 0.272 0.318 0.275 0.319 0.323 0.339 0.370 0.338 0.285 0.334 0.341 0.338 0.325 0.411 0.663 0.420 0.566 0.515
Solar MSE 0.192 0.230 0.243 0.243 0.227 2.58% 0.203 0.233 0.248 0.249 0.233 0.290 0.320 0.353 0.356 0.330 0.884 0.834 0.941 0.882 0.885 0.236 0.217 0.249 0.241 0.235
MAE 0.222 0.251 0.263 0.265 0.250 4.58% 0.237 0.261 0.273 0.275 0.262 0.378 0.398 0.415 0.413 0.401 0.711 0.692 0.723 0.717 0.711 0.259 0.269 0.283 0.317 0.280
RMSP 0.061 0.060 0.072 0.090 0.071 19.32% 0.080 0.089 0.090 0.093 0.088 0.363 0.374 0.389 0.385 0.378 0.449 0.563 0.694 0.868 0.644 0.059 0.075 0.105 0.150 0.097
MAPE 1.813 1.948 1.984 2.081 1.957 2.25% 1.842 2.052 1.986 2.129 2.002 2.225 2.378 2.558 2.633 2.449 2.349 2.403 2.509 2.042 2.326 2.200 2.471 2.775 2.531 2.494
sMAPE 0.381 0.426 0.437 0.443 0.422 32.15% 0.394 0.416 0.422 1.256 0.622 0.688 0.714 0.727 0.714 0.711 0.767 0.861 1.062 1.260 0.988 0.380 0.425 0.465 0.517 0.447
MASE 0.286 0.379 0.377 0.374 0.354 28.19% 0.301 0.369 0.362 0.939 0.493 0.436 0.511 0.519 0.507 0.493 0.562 0.670 0.817 0.905 0.739 0.273 0.344 0.363 0.415 0.349
Q25 0.225 0.242 0.254 0.258 0.245 32.13% 0.223 0.255 0.250 0.716 0.361 0.449 0.476 0.496 0.491 0.478 0.453 0.519 0.647 0.712 0.583 0.229 0.272 0.327 0.329 0.289
Q75 0.230 0.283 0.294 0.301 0.277 25.94% 0.250 0.253 0.277 0.715 0.374 0.308 0.321 0.336 0.336 0.325 0.504 0.544 0.636 0.682 0.592 0.229 0.262 0.280 0.333 0.276
ILI MSE 2.065 1.917 1.966 2.114 2.016 10.88% 2.329 2.189 2.217 2.314 2.262 3.052 2.804 2.829 2.973 2.915 3.410 3.365 3.125 2.847 3.187 5.421 5.001 5.098 5.312 5.208
MAE 0.835 0.852 0.854 0.900 0.860 10.23% 0.939 0.946 0.956 0.989 0.958 1.245 1.153 1.161 1.193 1.188 1.296 1.252 1.200 1.146 1.224 1.606 1.549 1.564 1.584 1.576
RMSP 0.341 0.371 0.371 0.386 0.367 11.57% 0.393 0.413 0.421 0.434 0.415 0.608 0.550 0.561 0.576 0.574 0.643 0.590 0.581 0.570 0.596 0.827 0.791 0.778 0.773 0.792
MAPE 3.701 3.193 3.054 2.874 3.206 -3.05% 3.638 3.169 2.937 2.701 3.111 4.334 2.953 2.662 2.504 3.113 5.282 4.880 4.463 3.753 4.595 2.545 2.625 2.545 2.260 2.494
sMAPE 0.633 0.657 0.649 0.660 0.650 7.54% 0.683 0.702 0.707 0.718 0.703 0.903 0.840 0.849 0.866 0.865 0.957 0.897 0.899 0.882 0.909 1.203 1.175 1.172 1.181 1.183
MASE 0.557 0.520 0.549 0.668 0.574 11.42% 0.622 0.587 0.629 0.753 0.648 0.860 0.715 0.726 0.861 0.791 0.777 0.718 0.716 0.747 0.740 0.932 0.856 0.861 0.980 0.907
Q25 0.821 0.761 0.742 0.740 0.766 5.20% 0.873 0.798 0.789 0.771 0.808 0.986 0.825 0.804 0.792 0.852 1.248 1.150 1.054 0.941 1.098 0.824 0.833 0.857 0.876 0.848
Q75 0.849 0.943 0.966 1.061 0.955 13.73% 1.005 1.094 1.123 1.207 1.107 1.503 1.480 1.518 1.593 1.524 1.345 1.354 1.346 1.350 1.349 2.388 2.265 2.271 2.293 2.304
OiK_Weather MSE 0.631 0.686 0.706 0.733 0.689 4.31% 0.672 0.718 0.734 0.755 0.720 0.662 0.697 0.713 0.731 0.701 0.733 0.820 0.899 0.960 0.853 0.569 0.647 0.713 0.752 0.670
MAE 0.582 0.616 0.630 0.645 0.618 2.22% 0.601 0.630 0.642 0.654 0.632 0.613 0.635 0.645 0.656 0.637 0.642 0.696 0.739 0.769 0.712 0.560 0.617 0.660 0.662 0.625
RMSP 0.712 0.770 0.790 0.820 0.773 1.02% 0.727 0.776 0.800 0.820 0.781 0.764 0.800 0.817 0.837 0.805 0.777 0.865 0.925 0.966 0.883 0.712 0.787 0.855 0.825 0.795
MAPE 4.005 4.934 5.270 4.378 4.647 41.81% 9.158 7.506 7.768 7.510 7.986 5.019 4.616 4.613 4.400 4.662 10.976 8.417 9.650 9.859 9.726 6.597 7.261 7.739 7.372 7.242
sMAPE 1.024 1.079 1.101 1.125 1.082 0.73% 1.040 1.085 1.110 1.126 1.090 1.133 1.178 1.200 1.229 1.185 1.087 1.175 1.243 1.299 1.201 1.035 1.123 1.201 1.143 1.126
MASE 0.897 0.888 0.889 0.876 0.888 5.03% 0.947 0.921 0.939 0.931 0.935 0.966 0.944 0.935 0.925 0.943 0.878 0.943 1.042 1.103 0.992 0.781 0.814 0.868 0.807 0.818
Q25 0.563 0.605 0.629 0.638 0.609 3.03% 0.602 0.624 0.633 0.652 0.628 0.609 0.631 0.641 0.652 0.633 0.668 0.702 0.739 0.768 0.719 0.547 0.608 0.642 0.645 0.611
Q75 0.600 0.628 0.631 0.653 0.628 1.26% 0.600 0.636 0.651 0.655 0.636 0.616 0.639 0.649 0.661 0.641 0.617 0.690 0.739 0.771 0.704 0.574 0.626 0.679 0.680 0.640
NN5 MSE 0.798 0.737 0.680 0.659 0.719 0.55% 0.793 0.745 0.685 0.668 0.723 1.474 1.484 1.406 1.429 1.448 0.891 0.915 0.739 0.859 0.851 0.969 0.983 0.939 0.977 0.967
MAE 0.590 0.582 0.570 0.568 0.578 0.86% 0.593 0.589 0.574 0.574 0.583 0.929 0.943 0.916 0.927 0.929 0.659 0.683 0.617 0.678 0.659 0.714 0.734 0.725 0.733 0.727
RMSP 0.526 0.539 0.545 0.554 0.541 2.35% 0.536 0.554 0.557 0.570 0.554 0.942 0.988 0.964 0.980 0.969 0.614 0.653 0.606 0.669 0.636 0.702 0.750 0.749 0.753 0.739
MAPE 2.332 2.563 2.426 2.431 2.438 5.28% 2.491 2.640 2.560 2.605 2.574 3.845 3.785 3.848 3.763 3.810 2.700 2.936 2.642 2.760 2.760 2.698 2.893 2.823 3.006 2.855
sMAPE 0.848 0.862 0.863 0.871 0.861 1.82% 0.860 0.877 0.878 0.892 0.877 1.276 1.326 1.296 1.316 1.304 0.954 0.992 0.941 1.014 0.975 1.046 1.092 1.096 1.088 1.081
MASE 0.528 0.541 0.530 0.519 0.530 2.21% 0.538 0.553 0.543 0.534 0.542 0.820 0.856 0.830 0.833 0.835 0.589 0.623 0.572 0.628 0.603 0.626 0.685 0.670 0.667 0.662
Q25 0.598 0.607 0.594 0.587 0.597 0.67% 0.601 0.611 0.597 0.593 0.601 0.876 0.897 0.867 0.880 0.880 0.679 0.708 0.644 0.698 0.682 0.730 0.730 0.748 0.717 0.731
Q75 0.582 0.558 0.546 0.550 0.559 1.06% 0.585 0.568 0.550 0.555 0.565 0.982 0.989 0.966 0.974 0.978 0.640 0.659 0.590 0.659 0.637 0.698 0.739 0.702 0.748 0.722
Rideshare MSE 0.225 0.261 0.467 0.504 0.364 26.61% 0.247 0.273 0.580 0.885 0.496 0.933 1.148 1.063 0.967 1.028 0.339 0.514 0.791 0.793 0.609 0.846 0.435 0.537 0.628 0.612
MAE 0.273 0.296 0.412 0.445 0.357 13.77% 0.291 0.300 0.473 0.591 0.414 0.771 0.864 0.824 0.801 0.815 0.416 0.528 0.673 0.681 0.575 0.628 0.399 0.472 0.524 0.506
RMSP 0.177 0.189 0.256 0.310 0.233 11.74% 0.187 0.186 0.312 0.369 0.264 0.777 0.874 0.866 0.842 0.840 0.401 0.503 0.632 0.638 0.544 0.494 0.248 0.332 0.384 0.365
sMAPE 0.388 0.406 0.589 0.652 0.509 11.48% 0.420 0.423 0.675 0.781 0.575 1.220 1.314 1.288 1.309 1.283 0.603 0.817 1.022 1.038 0.870 0.865 0.554 0.657 0.721 0.699
MASE 1.401 1.294 1.218 1.275 1.297 -3.10% 1.342 1.299 1.183 1.208 1.258 1.025 1.030 1.047 1.094 1.049 1.074 1.060 1.066 1.128 1.082 1.027 1.052 1.073 1.137 1.072
Q25 0.202 0.230 0.247 0.258 0.234 8.24% 0.217 0.211 0.272 0.320 0.255 0.417 0.459 0.437 0.421 0.434 0.330 0.300 0.355 0.351 0.334 0.355 0.257 0.285 0.318 0.304
Q75 0.344 0.362 0.578 0.633 0.479 16.40% 0.365 0.390 0.675 0.863 0.573 1.126 1.270 1.211 1.182 1.197 0.503 0.757 0.992 1.011 0.816 0.902 0.541 0.658 0.730 0.708
Univariate DeepBooTS iTransformer N-Beats N-Hits Autoformer
Dataset Metric Y1 Y2 Y3 Y4 Avg IMP Y1 Y2 Y3 Y4 Avg Y1 Y2 Y3 Y4 Avg Y1 Y2 Y3 Y4 Avg Y1 Y2 Y3 Y4 Avg
M4 Hourly MSE 0.177 0.203 0.242 0.270 0.223 28.06% 0.277 0.290 0.326 0.345 0.310 0.326 0.335 0.342 0.345 0.337 0.300 0.342 0.372 0.37 0.346 0.598 0.653 0.587 0.703 0.635
MAE 0.245 0.270 0.304 0.328 0.287 26.41% 0.369 0.373 0.401 0.415 0.390 0.412 0.423 0.427 0.431 0.423 0.39 0.421 0.433 0.428 0.418 0.583 0.621 0.589 0.649 0.611
RMDSPE 0.182 0.215 0.248 0.275 0.230 38.83% 0.354 0.354 0.389 0.405 0.376 0.375 0.394 0.396 0.403 0.392 0.367 0.41 0.418 0.414 0.402 0.576 0.635 0.592 0.651 0.614
MAPE 1.364 1.478 1.688 1.819 1.587 27.43% 2.108 2.109 2.224 2.305 2.187 2.532 2.576 2.510 2.614 2.558 2.344 2.523 2.569 2.531 2.492 3.830 4.019 3.701 4.369 3.980
sMAPE 0.445 0.488 0.534 0.569 0.509 22.41% 0.631 0.636 0.669 0.687 0.656 0.680 0.701 0.707 0.711 0.700 0.644 0.689 0.708 0.705 0.687 0.853 0.896 0.865 0.909 0.881
MASE 0.226 0.265 0.292 0.322 0.276 24.59% 0.339 0.354 0.376 0.395 0.366 0.365 0.396 0.399 0.398 0.390 0.359 0.395 0.415 0.415 0.396 0.481 0.522 0.521 0.546 0.518
Q25 0.230 0.249 0.276 0.294 0.262 29.00% 0.353 0.354 0.379 0.389 0.369 0.344 0.351 0.358 0.360 0.353 0.38 0.4 0.407 0.394 0.395 0.460 0.451 0.449 0.563 0.481
Q75 0.259 0.291 0.332 0.362 0.311 24.15% 0.384 0.392 0.424 0.441 0.410 0.479 0.496 0.496 0.502 0.493 0.401 0.442 0.46 0.463 0.442 0.707 0.792 0.729 0.735 0.741
Us_births MSE 0.316 0.401 0.334 0.405 0.364 3.70% 0.353 0.437 0.354 0.367 0.378 0.562 0.705 0.608 0.804 0.670 0.371 0.528 0.448 0.806 0.538 0.545 0.812 0.984 1.264 0.901
MAE 0.391 0.459 0.412 0.462 0.431 1.82% 0.413 0.473 0.434 0.436 0.439 0.586 0.653 0.599 0.715 0.638 0.467 0.579 0.529 0.737 0.578 0.583 0.709 0.793 0.943 0.757
RMDSPE 0.217 0.267 0.235 0.271 0.248 1.20% 0.228 0.268 0.252 0.256 0.251 0.376 0.414 0.377 0.469 0.409 0.283 0.363 0.331 0.477 0.364 0.356 0.417 0.458 0.558 0.447
MAPE 0.754 0.765 0.709 0.784 0.753 3.71% 0.776 0.817 0.813 0.720 0.782 0.974 1.121 0.932 1.037 1.016 0.851 1.024 0.879 1.201 0.989 0.844 0.968 1.235 1.447 1.124
sMAPE 0.403 0.462 0.416 0.497 0.445 -1.60% 0.409 0.466 0.420 0.458 0.438 0.585 0.628 0.590 0.698 0.625 0.447 0.536 0.492 0.684 0.540 0.634 0.762 0.819 0.940 0.789
MASE 0.328 0.393 0.351 0.381 0.363 0.55% 0.336 0.380 0.376 0.367 0.365 0.504 0.560 0.513 0.605 0.546 0.390 0.486 0.438 0.542 0.464 0.546 0.676 0.768 0.956 0.737
Q25 0.462 0.517 0.438 0.590 0.502 -2.45% 0.497 0.559 0.395 0.510 0.490 0.440 0.459 0.421 0.450 0.443 0.309 0.356 0.332 0.419 0.354 0.587 0.676 0.753 0.891 0.727
Q75 0.321 0.401 0.387 0.335 0.361 6.96% 0.329 0.386 0.472 0.363 0.388 0.732 0.846 0.777 0.979 0.834 0.624 0.801 0.727 1.056 0.802 0.580 0.743 0.833 0.995 0.788
Sunspot MSE 1.185 1.265 1.170 1.155 1.194 1.00% 1.207 1.286 1.175 1.156 1.206 1.044 1.049 1.009 1.010 1.028 0.990 1.004 0.971 0.961 0.982 1.220 1.290 1.224 1.275 1.252
MAE 0.535 0.564 0.542 0.536 0.544 1.45% 0.542 0.572 0.548 0.545 0.552 0.539 0.550 0.541 0.540 0.543 0.544 0.532 0.526 0.525 0.532 0.607 0.660 0.640 0.652 0.640
RMDSPE 0.805 0.838 0.798 0.789 0.808 2.88% 0.810 0.868 0.822 0.828 0.832 0.960 0.984 0.954 0.949 0.962 0.880 0.887 0.823 0.870 0.865 1.024 1.087 1.046 1.073 1.058
MAPE 2.725 2.716 2.818 2.844 2.776 2.87% 2.817 2.824 2.886 2.906 2.858 1.896 1.941 2.090 2.142 2.017 2.770 2.315 2.729 2.575 2.597 2.487 2.928 2.775 3.033 2.806
sMAPE 1.044 1.059 1.056 1.051 1.053 1.22% 1.028 1.083 1.070 1.083 1.066 1.406 1.419 1.369 1.352 1.387 1.186 1.214 1.139 1.216 1.189 1.349 1.416 1.396 1.317 1.370
MASE 1.413 1.315 1.134 1.065 1.232 2.07% 1.518 1.275 1.133 1.107 1.258 1.069 0.925 0.891 0.885 0.943 0.926 0.863 0.811 0.830 0.858 1.022 1.014 1.060 1.130 1.057
Q25 0.391 0.397 0.416 0.416 0.405 2.64% 0.391 0.406 0.426 0.442 0.416 0.509 0.514 0.514 0.518 0.514 0.523 0.486 0.488 0.512 0.502 0.545 0.616 0.619 0.641 0.605
Q75 0.669 0.701 0.670 0.657 0.674 1.89% 0.692 0.738 0.670 0.648 0.687 0.569 0.586 0.568 0.563 0.572 0.566 0.578 0.565 0.538 0.562 0.669 0.704 0.660 0.664 0.674
Saugeenday MSE 0.276 0.31 0.362 0.56 0.377 1.31% 0.279 0.312 0.366 0.569 0.382 0.276 0.308 0.357 0.512 0.363 0.282 0.312 0.362 0.505 0.365 0.296 0.324 0.373 0.607 0.400
MAE 0.371 0.397 0.431 0.543 0.436 0.91% 0.374 0.399 0.436 0.551 0.440 0.378 0.405 0.444 0.552 0.445 0.384 0.401 0.449 0.534 0.442 0.393 0.412 0.442 0.584 0.458
RMDSPE 0.367 0.398 0.439 0.575 0.445 1.11% 0.369 0.4 0.444 0.587 0.450 0.373 0.406 0.458 0.657 0.474 0.382 0.404 0.456 0.608 0.463 0.392 0.413 0.453 0.636 0.474
MAPE 71.15 78.47 86.28 100.96 84.21 1.36% 71.77 78.79 87.96 102.97 85.37 67.265 71.937 77.889 74.39 72.87 63.643 70.235 81.053 77.865 73.20 72.197 78.669 87.303 107.281 86.36
sMAPE 0.671 0.709 0.752 0.892 0.756 0.66% 0.674 0.711 0.758 0.901 0.761 0.687 0.732 0.793 1.003 0.804 0.701 0.724 0.791 0.949 0.791 0.703 0.73 0.768 0.944 0.786
MASE 1.211 1.234 1.323 1.688 1.364 1.23% 1.229 1.293 1.34 1.66 1.381 1.199 1.235 1.307 1.357 1.275 1.282 1.268 1.367 1.434 1.338 1.092 1.143 1.229 1.556 1.255
Q25 0.369 0.397 0.425 0.536 0.432 2.92% 0.375 0.401 0.443 0.561 0.445 0.387 0.42 0.465 0.584 0.464 0.394 0.393 0.49 0.537 0.454 0.406 0.431 0.447 0.639 0.481
Q75 0.374 0.393 0.435 0.535 0.434 0.23% 0.373 0.397 0.429 0.542 0.435 0.368 0.391 0.422 0.519 0.425 0.373 0.408 0.409 0.53 0.430 0.38 0.394 0.437 0.53 0.435
Appendix S Visualization of TS forecasting
For clarity and comparison among different models, we present supplementary prediction showcases for three representative datasets in Fig. 15, 16, and 17. Visualization of different models for qualitative comparisons. Prediction cases from the Traffic, Electricity and Weather datasets. These showcases correspond to predictions made by the following models: iTransformer (liu2023itransformer) and PatchTST (nie2022time), DLinear (zeng2023transformers), Autoformer (wu2021autoformer) and Informer (Zhou2021Informer). Among the various models considered, the proposed DeepBooTS stands out for its ability to predict future series variations with exceptional precision, demonstrating superior performance.
Figure 15: Prediction cases from the Traffic dataset under the input-96-predict-96 setting.
Figure 16: Prediction cases from the Electricity dataset under the input-96-predict-96 setting.
Figure 17: Prediction cases from the Weather dataset under the input-96-predict-96 setting.